### v1

# Estimating Spatially Resolved Radiation Fields Using Neural Networks

Abstract
We present an in-depth analysis on how to build and train neural networks to estimate the spatial distribution of scattered radiation fields for radiation protection dosimetry in medical radiation fields, such as those found in Interventional Radiology and Cardiology. Therefore, we present three different synthetically generated datasets with increasing complexity for training, using a Monte-Carlo Simulation application based on Geant4. On those datasets, we evaluate convolutional and fully connected architectures of neural networks to demonstrate which design decisions work well for reconstructing the fluence and spectra distributions over the spatial domain of such radiation fields. All used datasets as well as our training pipeline are published as open source in separate repositories.
Code: https://github.com/Centrasis/radfield3d-nn
keywords:
Interventional Radiology, Monte-Carlo Simulation, FCNN, U-Net, Dataset, X-ray
articletype: Paper
OSL
Optically Stimulated Luminescence
TLD
Thermoluminescence Dosimeter
APD
Active Personal Dosimeter
EPD
Electronic Personal Dosimeter
MCS
Monte-Carlo Simulation
DSR
Dose-Structure Report
CT
Computed Tomography
MRI
Magnetic Resonance Imaging
CG
Computer Graphics
CNN
Convolutional Neural Network
PRT
Precomputed Radiance Transfer
NeRF
Neural Radiance Field
MLP
Multilayer Perceptron
CUDA
Compute Unified Device Architecture
PACS
Picture Archiving and Communication System
IR
Interventional Radiology
VR
Virtual Reality
AR
Augmented Reality
API
Application Programming Interface
FCNN
Fully Connected Neural Network
CNN
Convolutional Neural Network
RAF
Realistic Anthropomorphic Flexible
NVS
Novel View Synthesis
LSTM
Long-Short-Term-Memory
RNN
Recurrent Neural Network
SMAPE
Symmetric Mean Absolute Percentage Error
GPR
Gamma Passing Ratio
IoU
Intersection over Union
SSIM
Structural Similarity Index Measure
MADD
Maximum Allowed Dose Difference
FiLM
Feature-wise Linear Modulation
GMU
Gated Multimodal Unit
ICG
Institute of Computer Graphics of the Technical University Braunschweig
PTB
Physikalisch-Technische Bundesanstalt
ECAP
Erlangen Centre for Astroparticle Physics
SCK CEN
Belgian Nuclear Research Centre
WP
Work Package
PBRFNet
Parametric Beam Radiation Field Network
SPERFNet
Spectral Enhanced Radiation Field Network
SRBFNet
Static Rotatable Beam Field Network
1 Introduction
The spatial distribution of radiation in a volume is of interest for various applications in radiation protection and occupational dosimetry. Knowledge gained from the analysis of irradiation scenarios can be used for optimizing radiation exposure. This is important for work in nuclear power plants, radioactive waste disposal, or in radiological medicine, as Interventional Radiology (IR), in which medical staff performs surgeries under fluoroscopy. The current state of the art relies on the utilization of personal dosemeters, which are sufficient for spatially homogeneous radiation distributions, but inherently misestimate personal doses in inhomogeneous radiation fields.
In the context of IR, medical personnel are exposed to inhomogeneous radiation fields due to their proximity to the patient, which complicates the accurate assessment of individual doses. The reliability of current personal dosimetry methods is contingent upon the uniform distribution of radiation fields. However, this uniformity assumption does not hold in IR settings. To address this issue, previous research has proposed the use of computational dosimetry systems [1], which are designed to monitor the locations and postures of all individuals involved in an IR procedure. In order to assess doses in these exposure scenarios, such systems are making use of currently available methods, that are mainly based on Monte-Carlo Simulations [2].
Nonetheless, the existing MCSs of radiation transport lack the necessary speed for real-time dose calculations, a limitation that also affects the software used for Virtual Reality (VR) or Augmented Reality (AR) training. VR training systems can play a pivotal role in diminishing radiation exposure during interventions by enhancing the radiation awareness of medical staff, as evidenced by at least two research projects [3, 4]. Notwithstanding their significance, there is a paucity of reliable and realistic radiation dosimetry data to be used in real-time visualization with these systems.
To advance radiation protection, it is necessary to develop accelerated methods for the simulation of radiation transport. In this research work, we explore the surrogate model approach for MCS acceleration, whereby we learn spatially resolved radiation field distributions from a Monte-Carlo Simulation. Subsequently, the neural networks serve as efficient models for radiation-field reconstruction, providing a real-time capable radiation-transport simulation method. Furthermore, we have generated and made available a variety of datasets for the training of our models and analogous ones. The generation of our datasets was facilitated by the extension and utilization of the RadField3D [5] simulation software which itself is build upon the Geant4 [6] MCS framework to ensure physical correctness, as Geant4 is well established and validated for the targeted use-case.
Our primary innovation is the creation of three datasets of spatially resolved, three-dimensional fluence and spectra distributions for the training and evaluation of neural networks for the reconstruction of radiations fields for IR procedures. We accompany those with concrete neural network implementations and design recommendations based on ablation studies. Additionally, we define several metrics useful for evaluating and comparing spatially resolved radiation field estimators.
1.1 Monte-Carlo Simulation in Radiation Protection
The technique of using MCS to calculate the radiation transport of complex measurements is already widely spread in various disciplines of dosimetry, from radiation protection to medical physics. Therefore, a set of well-established and tested general-purpose MCS toolkits for radiation transport already exists, such as Geant4 [6], MCNP [7], and EGSnrc [8]. Since these toolkits are meant for general purpose applications ranging from astrophysics over nuclear reactor physics to dosimetry calculations, fine-tuning the MCS parameters for each specific use case is required. There are frameworks built upon these toolkits, such as GATE [9], which is based on Geant4, and it has already been used for assessing doses to medical staff in nuclear medicine by tailoring simulation models on the basis of real tracking of personnel movements [10]. However, as highlighted by Rondon et al., current MCS capabilities are still far from real-time as they require, at best, several tens of seconds if not minutes to be performed. This is particularly challenging in IR, as the calculation of radiation-fluence distributions in the space around the patient during interventions is computationally demanding and can require hours of simulations. However, real-time simulations capabilities are becoming more and more necessary for computational radiation protection dosimetry of staff, as proposed by prior research projects, such as the PODIUM project [1].
Besides the real-time dosimetry, the AR and VR applications for training interventional radiologists and associated personnel demand even faster execution speeds. Nevertheless, especially in AR and VR we can trade absolute accuracy for speed within reasonable bounds. Those applications were already presented various times of the last years, for example in 2012 by an c-arm training extension for VirtX [11] or by the perceptual study of Rainford et al.in 2023 [3]. All of the already implemented systems would benefit from physically correct radiation fields to display during training, but currently those systems are only able to render radiation fields generated by simplified deterministic algorithms due to strict timing requirements. For those applications, achieving frame rates of at least $90-120\,\mathrm{fps}$ requires total rendering times per volume of approximately $8-11\,\mathrm{ms}$ to avoid perceptible visual latency [12].
1.2 Computational Personal Dosimetry
Real-world measurements in IR scenarios employing Active Personal Dosimeters have already revealed the complexity of contemporary radiation fields and high local dose rates that professionals are exposed to [13]. Those measurements also implied that the definition of a representational location for wearing a dosimeter on the body of the medical staff (whether it is an APD or legal passive dosimeter) is not possible, and therefore, it is likely to both under- or overestimate the received doses by the staff. In the course of the PODIUM project, the researchers developed a camera-based tracking system that was integrated with a MCS application based on the MCNP [7] toolkit. The simulation divided the room into frustums projected from the surface of a sphere placed around the patient at the isocenter of the scene [14, 15]. The evaluation of each event was conducted as it traversed one of the surface segments of the sphere, effectively projecting it along the normal of that surface segment through the scene. The validation of the simulation was based on these frustums, with measurements of APDs worn by medical personnel during real interventions serving as the ground-truth. These interventions involved tracking the locations and postures of staff members.
1.3 Hybrid Methods and Surrogate Models for Radiation Transport MCS
Due to its notable slowness, as evidenced in the research by Vanhavere & Van Hoey [16], the MCS is ill-suited for real-time or even interactive applications. To address this issue, recent studies have explored the prospect of enhancing its calculation speed. These efforts entail the replacement of the most computationally inefficient components of the radiation transport simulations, which is refereed to as a hybrid approach. Those hybrid methods can be composed of denoising algorithms. The most promising results in this field are delivered by deep learning methods, that use artificial neural networks for transforming the noisy results of an early aborted MCS to those of a fully executed MCS, but with shorter execution times [17, 18, 19]. Although this approach has been the subject of more extensive research, it exhibits a significant disadvantage, as the MCS needs a certain minimum of statistics to be denoise-able which inevitably comes at the cost of a lower processing speed compared to a single neural network. Consequently, it does not meet the requirements for real-time applications especially in the context of VR and AR applications that we mentioned before.
Other approaches focus on the complete replacement of the MCS algorithm, with artificial neural networks. This innovative approach is called surrogate models and its feasibility was already demonstrated in a few studies for different scenarios [20, 21, 22, 23]. Most of these studies were conducted in the context of nuclear power plants, thus those methods aim to predict neutron radiation fields induced by nuclear isotopes. Therefore, those works were heavily focused on modeling shielding and its positioning regarding a static radiation source. For this use case, Wang et al.proposed the use of a residual Fully Connected Neural Network (FCNN) for predicting neutron radiation fields under various shielding configurations [23]. As they could assume a static radiation source, they split the room into regions with similar conditions regarding the radiation field and trained multiple networks for each region. Concurrently, Ye et al.implemented a 3D- Convolutional Neural Network (CNN) for 3D reconstruction in a similar scenario [21]. This 3D-CNN takes the unshielded radiation fields and shielding voxel geometry, convolutes those two channels down to a flat vector which is then rearranged to a 3D radiation field after the shielding process. Another use case for FCNN is demonstrated in the research work of Yisheng et al.[22], where a FCNN is used to predict the spatially resolved photon flux in a room based on the data of a single detector at a known position. The output of the neural network is a low resolution grid with $2000$ voxels. An example of the use of neural networks in the domain of this research work in the context of IR was provided by the research work of M. Villa Arias [20]. Here, a 3D-UNet is used to transform a CT-Scout-Scan to the surrounding radiation field around the patient induced by the CT-scanner.
2 Materials and Methods
As there is a lack of public available and in this context usable datasets, our first contribution is the generation and provision of three datasets for training spatially resolved doserate predicting neural networks. For those three datasets, we implemented two distinct neural networks architectures, with additional variants for each, in order to analyze the influence of those architectures and specific hyperparameters on the quality of the predictions. For this purpose, we considered two basic architectures: A FCNN architecture inspired by Neural Radiance Fields [24] and a 3D-U-Net [25] architecture which is based on the CNN architecture, but with an extra transposed path for reconstructing all voxels at once.
2.1 Datasets
We generated and provided three synthetic datasets of radiation fields for different situations with increasing parameter spaces. The datasets were all generated using the RadField3D [5] dataset generator. The resulting datasets, consists of one radiation field per fixed set of parameters. Those fields are stored in the RadFiled3D file format and contain three channels. One for the photons of the direct X-ray beam, one for the resulting scatter fields and one containing the binary voxel occupation of the modeled geometry. Each of the first two channels contains the following information per voxel: The simulated volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$, the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ and the statistical error $\epsilon_{rel}$ as described in the research work about the MCS application RadField3D. Additionally, all input parameters, such as the X-ray tubes output spectra or the tubes position and direction, are stored per field.
DS-01:
The first dataset we generated, contains $1250$ radiation fields. Each field has a resolution of 50×50×50 voxels with each voxel having an extent of $2\,\mathrm{cm}$ in every dimension resulting in a total field extent of $1\,\mathrm{m}$ per dimension. The varying parameter in this first dataset is the direction of the beam with $\phi,\theta\in[0,2\pi)$, while the X-ray tubes output is fixed to a radiation quality of H-100 according to ISO 4037-1 [26] and the beam shape is fixed to a cone beam with an opening angle of $\angle_{\mathrm{Beam}}=10\degree$. The distance of the X-ray tube relative to the origin is $2.5\,\mathrm{m}$. At the origin of the field, the headless torso of an male alderson RANDO phantom is placed to act as the scatter object for the radiation beam. This setting is derived from the validation measurements conducted for the RadField3D data generator [5], and, although simplified, it is considered realistic for modeling this first exposure scenario.
DS-02:
The second dataset contains $2156$ radiation fields and extends the previous one by the radiation source spectra as an varying parameter. Here we used random, but realistic, spectra for modeling primary radiation in the context of IR. The range of the parameters was defined on the basis of empirical feedback from medical physicist and radiologists, mainly gained during the PODIUM project. Therefore, we defined an energy range for the tube output energy $E_{\mathrm{tube}}\in[40,125]\,\mathrm{keV}$. Further, we added aluminum with a random thickness $t_{\mathrm{Al}}\in[2.5,7.5]\,\mathrm{mm}$ and, optionally, copper, with a random-thickness $t_{\mathrm{Cu}}\in[0.0,0.9]\,\mathrm{mm}$ for simulating realistic filtering material used to improve the X-ray beam quality. As the anode material, we set tungsten, which is a standard reference material in IR imaging devices, and we used a random anode angle of $\angle_{\mathrm{Anode}}\in[8,12]\degree$. From that parameter space for the radiation source, we sampled sets of parameters and passed them to SpekPy [27] to generate reasonable X-ray spectra for the X-ray beams in this dataset.
DS-03:
In the third and most complex dataset, dynamic X-ray tube distancing is introduced. Therefore, a total of $3779$ distinct radiation fields was generated. For this dataset, we carried over the same base parameter space as described in the preceding datasets, and extended it by sampling the X-ray tube distance from the scenes isocenter as $d_{\mathrm{tube}}\in[35,75]\,\mathrm{cm}$. In order to make the dataset even more realistic in representing typical real-world scenarios, the X-ray beam was collimated to a rectangular shape, with fixed dimensions of $40\,\mathrm{cm}\times 30\,\mathrm{cm}$ in the plane passing through the origin.
In Tab. 1, we have summarized the distribution of each input parameter inside the three datasets and additionally the distribution of the contained fluences by providing the dynamic ranges $\mathrm{DR}_{\mathrm{dB}}$ and the Gini coefficient $G_{\mathrm{Gini}}$, which are defined as:
$\mathrm{DR}_{\mathrm{dB}}=10\log_{10}\Bigl(\frac{\max(\Phi^{3}_{\mathrm{\gamma}})}{\min(\Phi^{3}_{\mathrm{\gamma}})}\Bigr)$
$G_{\mathrm{Gini}}=\frac{1}{n^{2}\mu}\sum_{i=1}^{n}x_{i}(2i-n-1)$
where $\mathrm{DR}_{\mathrm{dB}}$ is used to describe the actual ranges needed for reconstruction and $G_{\mathrm{Gini}}$ expresses the deviation of the present fluence distribution from a uniform distribution. For the calculation of $G_{\mathrm{Gini}}$, the fluences are expected to be sorted in ascending order. A Gini coefficient of $G_{\mathrm{Gini}}=0$ indicates, that the dataset is uniformly distributed, while $G_{\mathrm{Gini}}=1$ indicates, that the dataset contains only one value or category.
Table 1: Statistics over the presented datasets including all varying input parameters.
Dataset $\overline{\angle}_{\mathrm{Beam}}\pm\sigma$ $\overline{\mathrm{dist}}_{tube}\pm\sigma$ $\overline{\mathrm{kV}}\pm\sigma$ $\overline{\mathrm{kV_{p}}}\pm\sigma$ $\mathrm{DR}_{\mathrm{dB}}$ $G_{\mathrm{Gini}}$
DS-01 $10.0\degree\pm 0.0$ $2.5\,\mathrm{m}\pm 0.0$ $80.7\,\mathrm{kV}\pm 0.0$ $100\,\mathrm{kV}\pm 0.0$ $43.9\,\mathrm{dB}$ $0.790$
DS-02 $10.0\degree\pm 0.0$ $2.5\,\mathrm{m}\pm 0.0$ $48.7\,\mathrm{kV}\pm 9.3$ $82.1\,\mathrm{kV}\pm 24.6$ $53.5\,\mathrm{dB}$ $0.872$
DS-03 - $0.6\,\mathrm{m}\pm 0.1$ $48.7\,\mathrm{kV}\pm 9.3$ $82.1\,\mathrm{kV}\pm 24.6$ $76.0\,\mathrm{dB}$ $0.991$
By observing the differences between our three datasets, one can see, that by reducing the volume of the direct beam inside the radiation field volume, the dynamic range $\mathrm{DR}_{\mathrm{dB}}$ is also drastically increased. In the case of DS-03 when smaller opening angles of the beam and smaller distances of the X-ray tube from the isocenter are used, particle fluences $\Phi^{3}_{\mathrm{\gamma}}$ span over nearly $8$ orders of magnitudes, instead of the spanning of about $5$ orders of magnitudes for the other datasets. Concurrently, the distribution of fluences is more challenging, as the amount of voxels with fluences below $0.1\,\%$ of the maximum fluence rises by a factor of approximately $253$ from $0.3\,\%$ to $76\,\%$.
2.2 Data normalization
Figure 1: Distribution of relative voxel fluences $\Phi^{3}_{\mathrm{\gamma}}$ inside datasets DS-01 and DS-02 starting from 0.0 on the X-axis.
Figure 2: Distribution of relative voxel fluences $\Phi^{3}_{\mathrm{\gamma}}$ inside datasets DS-01 and DS-02 starting from $10^{-1}$ on the X-axis.
Figure 3: Distribution of relative voxel fluences $\Phi^{3}_{\mathrm{\gamma}}$ inside dataset DS-03 starting from 0.0 on the X-axis.
Figure 4: Distribution of relative voxel fluences $\Phi^{3}_{\mathrm{\gamma}}$ inside dataset DS-03 starting from $10^{-1}$ on the X-axis.
In principle, radiation fields inside the dataset are already normalized, as the photon spectra are normalized to probability distributions $p(E_{\mathrm{\gamma}})$ and the fluences are normalized per single primary photon from the simulation. However, due to the numerical nature of the datasets, after loading a single radiation field from a dataset, we performed an extra normalization step for the data to limit the occurring gradients and avoid numerical instability as well as introducing bias because of high total values. Thus this extra normalization step effectively mitigates the severe class imbalance in all our dataset towards low fluence voxels. This has been visualized in the global fluence histograms in Fig. 4 for DS-01 and DS-02 and in Fig. 4 for DS-03. The histograms were generated using the MaxNorm function to transform the fluences to $[0,1]$. From both figures we can draw, that the majority of voxels has a fluence lower then $5\,\%$ of the maximum fluence, which makes sense, as the major part of the radiation field is indeed scatter radiation. However, this imbalance could potentially lead to a misreconstruction of the primary beam. Especially for DS-03 from Fig. 4, we can see that the class imbalance is also present inside the primary beam component. Not as severely expressed, datasets DS-01 and DS-02 exhibit an unequal distribution of fluences as well, as visible from Fig. 4. Therefore, we examine the influence of different normalization functions regarding the reconstruction capabilities, to find the best fit to reconstruct the high and low fluence regions with sufficient results.
$\mathrm{MaxNorm}(x)$ $=\frac{x}{\max(\mathbf{X})},$ $\quad x\in\mathbf{X}\subset\mathbb{R}^{+}$
$\Rightarrow\mathrm{MaxNorm}:\mathbb{R}^{+}\rightarrow[0,1]$
$\mathrm{MaxLogNorm}_{\alpha}(x)$ $=\frac{\ln(1+\alpha\,x)}{\max_{y\in\mathbf{X}}(\ln(1+\alpha\,y))},$ $\quad x\in\mathbf{X}\subset\mathbb{R}^{+}$
$\Rightarrow\mathrm{MaxLogNorm}:\mathbb{R}^{+}\rightarrow[0,1]$
The idea of using alternative normalizations in the interval of $[0,1]$ is to create a more equal distribution of gradient strengths over the range of values to reconstruct. We pose it as a hypothesis, that this is of importance for our use case, as the dosimetry is regarding relative instead of absolute errors per voxel. Thus, it can be beneficial to elevate differences between small fluences and flatten the differences between higher fluences to give small deviations on small scales more impact. A classical solution for this issue is to use a transformation to the log-space, which is in general a reasonable choice for the following training of neural networks. Unfortunately, due to numerical instabilities, the steep ascend of the logarithm at very small values can become unstable under inversion and thus break the bijective property required for normalization. This instability can be observed for fluences of a magnitude lower than $10^{-5}$. As it can be seen from Tab. 1, for DS-03 the interval of $10^{-7},10^{-8}]$ is of interest for the reconstruction as well. To mitigate this issue we added a scaling factor $\alpha$ prior to the log-transformation. We test two values for $\alpha$: $10^{3}$ and $10^{5}$. As an alternative way to encounter the high dynamic range, we doubled the range of our previously linearly spread out normalization from $\Phi^{3}_{\mathrm{\gamma}}\in[0,1]$ to $\Phi^{3}_{\mathrm{\gamma}}\in[-1,1]$ which additionally separates the low fluence voxels from the few high energy voxels by the sign. We distinguish between those two variants of the MaxNorm by indicating them as $\mathrm{MaxNorm}_{0}^{1}$ and $\mathrm{MaxNorm}_{-1}^{+1}$.
2.3 Architectures
In order to access the effectivity of our neural network architectures, we implemented each architectures neural network model in three variants to handle the different provided parameters for each dataset. With this approach, for each architecture there is one model that can only handle X-ray tube rotations, one that can additionally handle varying spectra and one that can additionally handle varying tube distances.
In general, our models are all estimating the relative topology of the spatial radiation field distribution conserving the relation of intensities between all voxels inside one specific radiation field. This is important, as the actual fluence scales between different settings of the beam would introduce a bias towards high X-ray tube currents and thus harm the training of the neural networks. On the downside, in the real world application, this inherently requires our approach to run concurrently to at least one area dosimeter that reports doserates at a known point in space. As those devices, especially those with real-time measurement capabilities, are known to exhibit a certain energy and thus spectrum dependence, one needs to correct those devices accordingly. This is specifically of importance in our targeted use case, as we aim to predict the scatter field around the patient, where the spectrum is not homogeneous and difficultly predictable. Thus, we decided to train all our networks not of the actual target measurand, e.g., air kerma rate $\dot{K}_{\mathrm{air}}$ or $\mathbf{H}^{*}(10)$, but on the components of those measurands as they are provided by RadField3D, which are the volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$ and the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ composed of $32$ bins and $E_{max}=150\,\mathrm{keV}$. In fact, the concurrent measurements of an area dosimeter are only needed to correct the fluence. Therefore, one can potentially use the estimated spectra to correct an area dosimeter measurement in turn.
2.3.1 Fully-Connected-Voxelwise-Networks.
The primary family of neural networks that we explored is the family of Fully Connected Neural Networks. Those networks are not only the most primitive, but also the most flexible ones. They have long been proven to be universal function approximators [28] and as such, given the right input representation, they can effectively learn to reconstruct structures of various kinds. More specifically, we were inspired by the NeRF approach of Mildenhall et al.[24] which is used for Novel View Synthesis (NVS). We chose this approach, as the reconstruction of radiance fields especially by follow up works for the real-time rendering application, seems promising for the reconstruction of radiation fields in real-time. As we chose to not only predict a scalar doserate per voxel, but the photon fluence $\Phi^{3}_{\mathrm{\gamma}}$ together with the normalized local photon spectrum $p(E_{\mathrm{\gamma}})$, this matches the NeRF architecture, where the reconstruction of visual density and color of a given point is separated as well.
For the training process of our FCNNs, we used a cosine annealing [29] learning rate scheduler with $\mathrm{eta}_{min}=10^{-6}$ together with an Adam optimizer using weight decay [30] of $10^{-4}$ and $\beta_{1}=0.9\land\beta_{2}=0.99$. Additionally, we preceded our main learning rate scheduler by a linear warmup phase that reaches our target learning rate after $1000$ steps.
Static Rotatable Beam Field Network (SRBFNet):
[…]
Figure 5: Schematic of the fully connected architecture used as the backbone for SRBFNet, SPERFNet and PBRFNet outputting a spectrum and a fluence for each voxel location.
This neural network forms the basis for reconstructing the presented datasets. For each dataset, we present variations of this network. In general, SRBFNet receives the same input parameters as the original NeRF did: The voxel location in cartesian coordinates ($x,y,z\in[0,1]$), normalized per dimension, and a normalized direction vector, which in this case is the X-ray tube direction. The overall structure is shown in Fig. 5, where global parameters g in this case is the encoded direction vector.
Our network begins by feeding the location vector through a frequency encoding using sine and cosine functions with different periodicities which was initially introduced by A. Vaswani et al.[31] for the transformer networks and reused for the original NeRF implementation. Even though, the direction vector could also be represented by a simple frequency encoding, we instead used a spherical harmonics based encoding as the cyclic nature of an rotation angle can better be represented by that encoding. This was demonstrated for NeRF-like networks by InstantNGP of T. Müller et al.[32]. For efficiency reasons, we use the spherical harmonics implementation as it is provided by the tiny-cuda-nn package [33] with version 2.0. For both encodings, we append the raw input to the encoded vector. In contrast to the classical NeRF application, our network does not aim to learn the scatter objects 3D geometry with visual densities for volumetric rendering. Thus, we skip the pre-processing path where the model should optimize the voxel location representation towards the local density independently from the direction. We want the network to learn the implicit representation of the fluence distribution, which is similar to the visual density, but in our case is ultimately depended from the radiation direction. Thus, we extract the fluence as well as the spectrum from the last layer. In order build a robust backbone for solving the radiation field estimation under more complex parameter spaces, we do not just concatenate the feature vectors that describe the global beam conditions like the view direction in NeRF. Instead, we experimented with various feature fusing methods and finally compared the Feature-wise Linear Modulation (FiLM) [34] approach, which was originally used in CNNs, with a Gated Multimodal Unit (GMU) [35] approach and as the baseline, with the simple concatenation of feature vectors like it is done in classical NeRF networks. Ahead of fusing both feature vectors, the encoded X-ray tube direction vector is passed through the global encoding block, a Multilayer Perceptron (MLP) block that is similar to the MLP processing the encoded location vector with one hidden layer and a model width that is equal to the width of the main path of the network.
The actual used hyperparameters for this and the extensions of this network, like frequencies of the frequency encoding, spherical harmonics degrees, model widths and feature fusing layers are reported in the results section as we determined them using hyperparameter tuning. Through all layers of the main path of this network, except for the last ones, we use SiLU activation functions. Only for the output activation we use different activation functions. For the fluence, we use a different activation function depending on the range of the fluence normalization. For the case of a normalization in $[0,1]$, we used a sigmoid $\sigma$ activation function, while for a normalization in [1,1][-1,1], we used a gradient conserving clamping $\mathrm{clamp}_{\mathrm{GC}}$, with:
$\sigma(x)=\frac{1}{1+\mathrm{exp}(-x)}$
${\mathrm{clamp}_{\mathrm{GC}}}_{min}^{max}(x)=x+(\mathrm{clamp}_{min}^{max}(x)-x)$
For the gradient conserving clamping, the second term of the sum is executed detached from pytorchs autograd graph which results in a clamping that still has the original gradients attached. For the local spectra, we use a simple histogram normalization $\mathrm{norm}_{\mathrm{hist}}$ as the activation function, to guarantee normalized spectra for the losses.
$\mathrm{norm}_{\mathrm{hist}}(x)=\frac{x}{\sum_{i=0}^{n-1}(x_{i})},n=32$
SRBFNet was designed to be trained on the parameter space provided by DS-01.
Spectral Enhanced Radiation Field Network (SPERFNet):
This network is the first extension of SRBFNet. It extends the global parameters g by the x-ray spectrum of the radiation source. Each radiation field file contains the $150$ bin histogram of the X-ray tube output spectrum using a bin width of $1\,\mathrm{keV}$. In order to improve the radiation field distribution and include the reconstruction of the local spectra, we encode that spectrum into our network. First, we reduce the dimensionality drastically by bilinear resampling the bins of the histogram down to $64$ bins. The reduced histogram is then passed through a MLP with one hidden layer using a width of $32$ neurons and the SiLU activation function. After the first linear layer, we implemented a layer-normalization to reshape the activation distribution.
The two feature vectors of the X-ray tube direction and the X-ray tube output spectrum are then concatenated and combined by the global encoding block in Fig. 5 that previously just processed the direction vector.
SPERFNet was designed to be trained on the parameter space provided by DS-02.
Parametric Beam Radiation Field Network (PBRFNet):
This neural network targets the limitation of the pervious described variant, that is only able to learn a fixed, static beam distance. Therefore, this model adds an additional encoder to allow for the additional beam parameter introduced by DS-03 while keeping the same overall structure from SPERFNet. This requires one new scalar input parameter, that is the tube distance as the opening angle of the rectangular beam is directly depending on the distance. We observed, that simply passing this scalar raw is leading the neural network to ignore it. The previously used frequency encoding is not useful here either, as the input data is neither of high frequency nor periodical. Thus, we added a mini MLPs into the architecture, with has a width of $16$ neurons and in an optimal way encodes the distance information for the global parameters vector g.
PBRFNet was designed to be trained on the parameter space provided by DS-03.
2.3.2 Convolutional-3DVolume-Networks.
Instead of predicting every single voxel by itself, which is inefficient when the aim is to always predict all voxels, especially for fine resolutions, one can use CNNs. Those networks are relying on using convolution kernels, with multiple channels, that sweep over the image space and by that encoding local information. Initially, those networks were only used for Computer Vision, but by transposing this operation, those networks can be used for generative purposes.
Beam2Scatter U-Net:
Our Beam2Scatter (B2S) U-Net is in general a CNN, that encodes the raw radiation field of the direct beam prior to the interaction with the patient geometry and from that should estimate the scatter radiation field as it is induced by the patient. Therefore, its encoder-decoder architecture follows the 3D-U-Net description [25], which itself is just an adaption to the third dimension of the original U-Net, as it was initially used for biomedical image segmentation [36]. We adopt this structure and give our network the fluence map of the direct beam as a one-channel input and let the network predict a $33$ dimensional output with the first $32$ channels for the spectra and the last channel for the fluence. Another addition is, that we introduced FiLM [34] layers after each up-sampling step in the decoder path, to modulate the weights of the decoder depending on the encoded input spectrum of the X-ray tube, which is encoded the same way as in SPERFNet. This way we can inject that information into the U-Net structure. The networks outputs got the prior introduced activation functions applied: ${\mathrm{clamp}_{\mathrm{GC}}}_{min}^{max}$ and $\mathrm{norm}_{\mathrm{hist}}$. Thanks to the direct beam radiation distribution being the input of this network, it can be directly used with all three datasets. For DS-01, the spectra encoding and fusing by using FiLM layers was disabled.
2.4 Training pipeline
To ensure that the models are all comparable, the training pipeline for all models was identical. Regardless, whether the model itself was capable of predicting partial or whole fields, the evaluation of the loss functions and the sub sequential update of the weights was performed in the same way. All models were trained for a maximum of $200$ epochs, which could be prematurely aborted, when the validation loss was not improving for the last $10$ epochs. For the training loop, we used the pytorch lightning [37] framework which implements the Cross-Validation policy with a final test using a dedicated validation and test dataset. The ratio of the split into training, validation and test dataset was $70\,\%$, $15\,\%$ and $15\,\%$. At first, a number of fields is loaded from the dataset that are transformed to pytorch tensors, preserving the channels and single layers, and as such they are batched together. We used a physical batch size of $4$ full radiation fields. Multiple batches of those predicted and ground-truth fields are aggregated together to an effective batch size of $64$ for training. We applied data pre-processings to each field prior to the prediction. Those pre-processings included the optional joining of the direct beam and scattered field channels, but also normalizing the ground-truth and input tensors. Finally, the loss functions for each of the neural networks normalized outputs (fluence and spectra) are calculated. Different from that, the metrics are calculated on the back-transformed data in the original data space.
The initial learning rates where estimated per neural network by using the pytorch lighting implementation of a learning rate finder, which follows the descriptions of L. N. Smith [38]. We limited it to search on the interval of $10^{-2},10^{-4}]$ for $250$ steps.
Per voxel training:
For the FCNN models, which operate per voxel, we added an optional extra step to the training pipeline, that is typical for NeRF approaches. We refer to it as voxel location randomization, which is essentially a perturbation of the voxel position around its center with a maximum jitter of half the voxels dimension. This was used as a data augmentation to stabilize training in NeRF variants like InstantNGP [32].
2.4.1 Loss functions.
Across all our models we used the same two loss functions, that we optimize concurrently: $\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}}$ the loss regarding the volumetric fluence $\Phi^{3}_{\mathrm{\gamma}}$ and $\pazocal{L}_{p(E_{\mathrm{\gamma}})}$ the loss regarding the local photon spectrum $p(E_{\mathrm{\gamma}})\in\mathbb{R}^{32}$ for $\sum_{i=0}^{31}p(E_{\mathrm{\gamma},i})=1$.
$\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}}(t,p)$ $=\frac{1}{3}\Bigl(\mathrm{L_{1}}(t,p)+\mathrm{L_{2}}(t,p)+\bigl(1-\mathrm{SSIM}(t,p)\bigr)\Bigr),\$$\mathrm{with}\ t,p\in\mathbb{R}$
[…]
$\mathrm{Wasserstein}(t,p)$ $=\frac{1}{n}\sum_{i=0}^{n-1}|\mathrm{CumSum}_{i}(p)-\mathrm{CumSum}_{i}(t)|,$ $\mathrm{with}\ t,p\in\mathbb{R}^{n}$
$\pazocal{L}_{p(E_{\mathrm{\gamma}})}(t,p)$ $=\frac{3}{10}L_{1}(t,p)+\frac{7}{10}\mathrm{Wasserstein}(t,p),$ $\mathrm{with}\ t,p\in\mathbb{R}^{32}$
With $t$ being the ground truth (target) and $p$ being the prediction of the neural network. Here, $L_{1}$ and $L_{2}$ refer to the L1- and L2-Norms using mean aggregation to produce a scalar loss for higher dimensional inputs.
2.4.2 Hyperparameter tuning.
For the hyperparameter tuning, we used the Optuna [39] framework with version 4.5. For the tuning of each network we defined the search space for possible parameters as listed in Tab. 2. The tuning was in general performed on every model we discuss: B2S U-Net, SRBFNet, SPERFNet, PBRFNet. Anyway, not all of the parameters could be tested for each model. For instance, the SRBFNet does not encode any spectra information, therefore the spectra encoding related parameters were not probed.
Table 2: Probed hyperparameter spaces for each parameter of the models (B2S U-Net, SRBFNet, SPERFNet, PBRFNet), where applicable.
Model parameter Parameter space
model width [64, 96, 128, 192, 256, 384]
location encoding frequencies [10, 12, 14]
direction encoding harmonics [4, 6, 8]
feature fusing [Concatenation, FiLM, ResFiLM, GMU]
normalizer $\mathrm{MaxNorm}_{0}^{1}$, $\mathrm{MaxNorm}_{-1}^{+1}$, $\mathrm{MaxLogNorm}_{10^{3}}$, $\mathrm{MaxLogNorm}_{10^{5}}$]
random voxel location [true, false]
encoded spectra dimensions [16, 32, 64]
2.5 Validation Metrics
To assess the quality of the developed models, we need to analyze their performance on the test section of our datasets. As the distribution of voxel values inside our dataset is special for our use case and we have special requirements as well, e.g., we do not aim for perceptual accuracy as in most other reconstruction tasks, we first need to assess which metrics and which variants of metrics are applicable for neural network evaluation of this kind. In general, we measure our metrics over the air kerma rate fields $\dot{K}_{\mathrm{air}}$ of ground-truth and predicted fields using the following three metrics classes.
Symmetric Mean Absolute Percentage Error (SMAPE):
The first and most obvious kind of metric we employed is the SMAPE which is providing the mean relative error across all voxels of the test dataset. As the dataset is strongly biased towards low voxel values we introduce this metric in two variants: $\mathrm{SMAPE}_{acc}^{x\,\%}$ for the metric applied over all voxels whose values exceed $\max(K_{air})*(1-\frac{x}{100})$ and $\mathrm{SMAPE}_{acc}^{scatter}$, which applies the metric on all voxels between $5\%$ and $0.5\%$ of $\max(K_{air})$.
The SMAPE is defined in the following, where $\mathbf{P}$ and $\mathbf{T}$ are two sets of voxels, present in two corresponding $K_{air}$ fields. $\mathbf{T}$ is the set of ground truth voxels and $\mathbf{P}$ is the set of predicted voxels:
$\mathrm{SMAPE}=\frac{2}{n}\sum_{i=1}^{n}{\frac{|P_{i}-T_{i}|}{|T_{i}|+|P_{i}|}}$
$\mathrm{SMAPE_{acc}}=1-\frac{1}{2}\mathrm{SMAPE}$
With $\mathrm{SMAPE_{acc}}$ referring to the SMAPE accuracy metric reported in this work.
Structural Similarity Index Measure (SSIM):
This metric is a well established approach of measuring the similarity in terms of structure between a target and a prediction image or volume [40]. This metric is commonly used for reconstruction tasks and expresses the structural similarity between target and prediction by an intuitive value between 0 and $1$, where $1$ indicates perfect accordance between both structures.
Gamma Passing Ratio (GPR):
The GPR within the scope of a gamma evaluation is a common metric in the field of radiotherapy radiation transport simulation validation [41]. It reflects the relative amount of voxels that pass a certain criterion within a given radiation field. The metric variant we implemented follows the idea of the Maximum Allowed Dose Difference (MADD) as described by Jiang et al.[42]. For our validation we used two different criteria. The strict criteria, which measures the amount of voxels with a MADD of $3\,\%$ using a spatial resolution of $4\,\mathrm{cm}$ and the more relaxed criterion, which measures the amount of voxels with a MADD of $10\,\%$ and a spatial resolution of $6\,\mathrm{cm}$. To calculate the GPR, the field is shifted by the allowed $\Delta_{d}$ in each dimension, respecting the extend of a single voxel. For $\Delta_{d}=4\,\mathrm{cm}$ this results in $2$ shifts per dimension for the case of a voxel size of $2\,\mathrm{cm}$. The minimum distance weighted error for each voxel across the applied shifts is used as the error for the scoring of the passing voxels, thus the error $\gamma(\vec{x})$ for a voxel in a voxel grid $\mathbf{X}$ at position $\vec{x}\in\mathbf{X}$ is defined as:
$\gamma^{\Delta_{d}}_{\Delta_{\dot{D}}}(\vec{x})=\min_{\vec{r}\in\mathbf{V}_{\Delta_{d}}(\vec{x})}\Biggl(\sqrt{\Bigl(\frac{\|\vec{x}-\vec{r}\|}{\Delta_{d}}\Bigr)^{2}+\Bigl(\frac{\dot{D}_{p}(\vec{x})-\dot{D}_{t}(\vec{r})}{\Delta_{\dot{D}}}\Bigr)^{2}}\Biggr),\mathbf{V}_{\Delta_{d}}(\vec{x})=\{\vec{v}\,|\,\|\vec{v}-\vec{x}\|<={\Delta_{d}}\land\vec{v}\in\mathbf{X}\}$
with $\dot{D}_{t}$ being the target doserate in $K_{air}$ at a given point and $\dot{D}_{p}$ being the predicted one.
The reported metric $\mathrm{GPR}_{p\,\%}^{d\,\mathrm{cm}}$ follows from that as:
$\mathrm{GPR}_{p\,\%}^{d\,\mathrm{cm}}=\frac{|\{\vec{x}\,|\,\gamma^{d}_{p}(\vec{x})\leq 1\land\vec{x}\in\mathbf{X}\}|}{|\mathbf{X}|}$
Spectrum Accuracy:
As the spectra represent a special kind of information with additional usage aside the direct dosimetry aspect, we evaluate them separately. Therefore, we use the Intersection over Union (IoU), where we calculate the ratio between the combined area of ground truth and prediction spectra and the overlapped area between them. This accuracy is defined as in the following, where $\mathbf{P}$ and $\mathbf{T}$ are two sets of voxels in two corresponding spectra radiation fields. $\mathbf{T}$ is the set of ground truth voxels and $\mathbf{P}$ is the set of predicted voxels:
$\mathrm{Spec}_{acc}=\frac{\sum_{i=1}^{n}{\min(P_{i},T_{i})}}{\sum_{i=1}^{n}{P_{i}+T_{i}-\min(P_{i},T_{i})}}$
3 Results
Table 3: Reported reconstruction metrics for each network in its optimal configuration of hyperparameters. All metrics were calculated from the test dataset section, grouped by used dataset.
Dataset Model $\mathrm{SMAPE}_{acc}^{90\,\%}$ $\mathrm{SMAPE}_{acc}^{scatter}$ SSIM $\mathrm{GPR}_{3\,\%}^{6\,\mathrm{cm}}$ $\mathrm{GPR}_{10\,\%}^{4\,\mathrm{cm}}$ $\mathrm{Spec}_{acc}$
DS-01 SRBF $\mathbf{96.4}\,\%$ $\mathbf{95.5}\,\%$ $\mathbf{0.961}$ $\mathbf{99.84\,\%}$ $\mathbf{99.98}\,\%$ $\mathbf{84.2}\,\%$
B2S U-Net $87.1\,\%$ $64.0\,\%$ $0.868$ $84.9\,\%$ $99.0\%$ $72.5\,\%$
DS-02 SRBF $90.0\,\%$ $83.3\,\%$ $0.923$ $6.8\,\% 99$$9.8\,\%$ $59.5\,\%$
SPERF $\mathbf{97.3}\,\%$ $\mathbf{96.2}\,\%$ $\mathbf{0.952}$ $\mathbf{99.7\,\%}$ $\mathbf{99.9}\,\%$ $\mathbf{86.2}\,\%$
B2S U-Net $88.0\,\%$ $55.5\,\%$ $0.844$ $72.7\,\%$ $98.1\%$ $77.1\,\%$
DS-03 PBRF $\mathbf{84.4}\,\%$ $\mathbf{84.8}\,\%$ $\mathbf{0.902}$ $\mathbf{87.2}\,\%$ $\mathbf{94.4}\,\%$ $\mathbf{86.7}\,\%$
B2S U-Net $19.5\,\%$ $19.4\,\%$ $0.802$ $21.8\,\%$ $30.5\,\%$ $14.2\,\%$
In general, the qualities of the developed models by this research work are condensed in Tab. 3. In there, each metric value belongs to a network, trained, tested and optimized in terms of hyperparameters on one of our datasets. The first dataset DS-01, was used to train our base FCNN architecture called SRBF. We compared the results of that network against the implementation of the U-Net that we introduced previously as Beam-to-Scatter (B2S) U-Net. As this network does not need any architectural changes to be applied on multiple datasets, as its inputs are the raw primary beam layer of a radiation field together with the X-ray tube output spectra, we used this as a reference for each dataset and variant of our FCNN architecture. We report the best metric values our networks achieved after we optimized the models hyperparameters by using Optuna as mentioned in the previous section. The specific found hyperparameters for each network, are listed in Tab. 4.
On DS-01, one can easily see, that the FCNN architecture outperforms the U-Net architecture. Even though, the U-Net also learns the structure of the radiation field, the edges between primary beam and scattered radiation are not reproduced as sharp as the FCNN does. This observation can be reproduced for DS-02. Additionally, we intended to highlight the effectiveness of our spectra encoding for the expected air kerma rates $\dot{K}_{\mathrm{air}}$. Thus, we trained the base SRBFNet model, which can not handle different output spectra of the X-ray tube, on DS-02. The effectivity of our spectra encoding can be seen by comparing SRBFNet against SPERFNet, where we get a nominal difference of $12.9\,\%$ for $\mathrm{SMAPE}_{acc}^{scatter}$ just from the missing spectra information. Furthermore, we can conclude, that training our FCNN architecture on a bigger dataset using more degrees of freedom, namely the X-ray tube spectrum, actually improves the $\mathrm{SMAPE}_{acc}$, even though the spatial resolution got slightly, but not notably, worse.
For DS-03 and PBRFNet, we observed, that using models with a higher neuron count by increasing the width of our model beyond $192$, e.g., $256$ or even $384$, the reconstruction capability is increasing by up to $1.5\,\%$ regarding $\mathrm{SMAPE}_{acc}^{scatter}$. A similar increase of the SSIM on the other hand side, and thus the overall shape of the radiation field, is invariant to a change of the model width beyond $192$ across all our datasets and models. As the benefit of using broader models is not significant enough to sacrifice performance for it, we limited the hyperparameter selection to models with a maximum model width of $192$.
Table 4: Used model configurations, where L denotes the number of frequencies used for the Fourier feature encoding of the voxel location, $l_{max}$ denotes the maximum degree used for the encoding of the radiation direction by spherical harmonics and $\mathrm{dim}(\mathbf{g}_{\mathrm{spec}})$ indicates the dimensionality of the encoded input spectrum. Model width refers to the number of neurons in one full layer of the MLP main path.
Dataset Model Model Width L $l_{max}$ $\mathrm{dim}(\mathbf{g}_{\mathrm{spec}})$ Norm Conditioning
DS-01 SRBF 192 10 4 - $\mathrm{MaxNorm}_{0}^{1}$ FiLM
B2S U-Net - - - - $\mathrm{MaxNorm}_{0}^{1}$ -
DS-02 SRBF 192 10 4 - $\mathrm{MaxNorm}_{0}^{1}$ FiLM
SPERF 192 10 4 16 $\mathrm{MaxNorm}_{0}^{1}$ FiLM
B2S U-Net - - - - $\mathrm{MaxNorm}_{0}^{1}$ FiLM
DS-03 PBRF 192 14 4 16 $\mathrm{MaxNorm}_{0}^{1}$ FiLM
B2S U-Net - - - - $\mathrm{MaxNorm}_{0}^{1}$ FiLM
Table 5: Inference durations per field ($50^{3}$ voxels) of each model, regarding a specific feature fusing method, using non-compiled pytorch models and a NVIDIA RTX 3090 GPU. All models were using the following configuration with a model width of $192$ and encoding parameters $l_{max}=4\land\mathrm{L}=10$ for SRBFNet and SPERFNet, with one exception for PBRFNet, which was using L = $14$.
Model Feature fusing Inference duration per field $\mathrm{SMAPE}_{acc}^{scatter}$
SRBF Concat $20.86\,\mathrm{ms}\ \pm\ 2.81\,\mathrm{ms}$ $94.3\,\%$
FiLM $20.55\,\mathrm{ms}\ \pm\ 1.87\,\mathrm{ms}$ $\mathbf{95.5}\,\%$
ResFiLM $24.46\,\mathrm{ms}\ \pm\ 1.82\,\mathrm{ms}$ $94.4\,\%$
GMU $41.00\,\mathrm{ms}\ \pm\ 2.55\,\mathrm{ms}$ $88.7\,\%$
SPERF Concat $21.76\,\mathrm{ms}\ \pm\ 2.57\,\mathrm{ms}$ $95.2\,\%$
FiLM $21.76\,\mathrm{ms}\ \pm\ 2.57\,\mathrm{ms}$ $\mathbf{96.2}\,\%$
ResFiLM $24.71\,\mathrm{ms}\ \pm\ 1.63\,\mathrm{ms}$ $95.4\,\%$
GMU $41.71\,\mathrm{ms}\ \pm\ 2.51\,\mathrm{ms}$ $93.7\,\%$
PBRF Concat $21.03\,\mathrm{ms}\ \pm\ 1.31\,\mathrm{ms}$ $79.2\,\%$
FiLM $22.17\,\mathrm{ms}\ \pm\ 2.54\,\mathrm{ms}$ $\mathbf{84.8}\,\%$
ResFiLM $24.32\,\mathrm{ms}\ \pm\ 1.88\,\mathrm{ms}$ $67.8\,\%$
GMU $40.43\,\mathrm{ms}\ \pm\ 2.42\,\mathrm{ms}$ $40.9\,\%$
4 Discussion
The configuration of the best performing models during the hyperparameter tuning are listed in Tab. 4. From those results, we can conclude, that for the fluence prediction, the same data range and activation function as used for NeRF-like approaches is beneficial, which is the use of sigmoid $\sigma$ together with a linear maximum normalization between $[0,1]$. Moreover, when additionally reviewing the table of Tab. 3, we are convinced, that using a model width of $192$ and a location encoding with frequencies within a range of $10$ to at max $14$ is sufficient for the presented use case. Consistently over all datasets, the angular resolution of the beam needed to solve this problem was found to be low enough, that a harmonics count of $4$ for the spherical harmonics encoding of the beam direction is sufficient. Further, two plain FiLM layers for feature fusing, one at an early and one at a late position in the network, was proven to be most effective for this use case compared to using GMU layers, Residual-FiLM layers and feature concatenation. Even though, the impact of the FiLM layers on the inference duration is measurable, it is small compared to the concatenation that is commonly used in the field of Novel View Synthesis, as it can be seen in Tab. 5. In that table, we concentrate on the differences regarding $\mathrm{SMAPE}_{acc}^{scatter}$, as this metric is reflecting the accuracy on the most interesting part of the predicted volume for the computational dosimetry. That is, because the staff during an IR procedure will be mainly monitored in the $\mathrm{SMAPE}_{acc}^{scatter}$ volume. From Tab. 5, one can also draw, that the use of FiLM layers did not only achieve the best accuracies, but also the shortest inference durations compared to GMU or Residual-FiLM. Moreover, we demonstrated, that the radiation transport simulation can be learned by the same deep, but narrow structure that is used to learn light transfer as in the use case of NVS. As the provided implementation does not use custom, native GPU codes like implementations in CUDA or OpenCL, its inference time does not enable real-time usage in the sense of the required $8\,\mathrm{ms}$ to $11\,\mathrm{ms}$ for VR- or AR-applications. Nevertheless, an inference time of about $20\,\mathrm{ms}$ is suitable for interactive applications. After all, with the provided FCNN models, state-of-the-art inferences times, that are suitable for the VR- and AR-application, should be achievable by implementing the network as CUDA or OpenCL kernels in the future, like it was already demonstrated by InstantNGP.
5 Conclusion
We have presented three datasets for training machine learning and especially deep learning algorithms on spatially resolved and physically accurate radiation field distributions as they could arise during IR procedures. Additionally, we have implemented and compared different variants of fully connected and convolutional neural networks, following the architectural design of U-Nets and NeRF-like architectures. Thereby, we have proven NeRF-like architectures to be efficient and capable of being used for radiation protection purposes and training of medical staff in the context of IR procedures. We have also pointed out, which design decisions are effective in this special context. Our presented datasets together with the presented findings on the architectural decisions for neural reconstruction networks create a profound foundation for enabling further research on this topic, e.g., extending our or similar approaches by uncertainty handling in neural networks or incorporation of dynamic scatter geometry.
Data availability
[…]
The authors gratefully acknowledge partial funding by the DFG under Germany’s Excellence Strategy within the Cluster of Excellence PhoenixD (EXC 2122, Project ID 390833453) and a DFG Research Grant (HU 2660/3-1, ”Development of real-time capable methods for the simulation of photon radiation - using the example of quantitative dosimetry in interventional radiology”, Project number 547148940).

### v2

@@
 We present an in-depth analysis on how to build and train neural networks to estimate the spatial distribution of scattered radiation fields for radiation protection dosimetry in medical radiation fields, such as those found in interventional radiology and cardiology.
-Therefore, we present three different synthetically generated datasets with increasing complexity for training, using a Monte-Carlo Simulation application based on Geant4.
+We present three different synthetically generated datasets with increasing complexity for training, using a Monte-Carlo Simulation application based on Geant4.
 On those datasets, we evaluate convolutional and fully connected architectures of neural networks to demonstrate which design decisions work well for reconstructing the fluence and spectra distributions over the spatial domain of such radiation fields.
-All used datasets as well as our training pipeline are published as open source in separate repositories.
+All our datasets, as well as our training pipeline, are published as open source in separate repositories.
 Code: https://github.com/Centrasis/radfield3d-nn
@@
 Knowledge gained from the analysis of irradiation scenarios can be used for optimizing radiation exposure.
-This is important for work in nuclear power plants, radioactive waste disposal, or in radiological medicine, as Interventional Radiology (IR), in which medical staff performs surgeries under fluoroscopy.
+This is important for work, for example, in nuclear power plants, radioactive waste disposal, or in radiological medicine, such as Interventional Radiology (IR), in which medical staff performs surgeries under fluoroscopy.
 The current state of the art relies on the utilization of personal dosemeters, which are sufficient for spatially homogeneous radiation distributions, but inherently misestimate personal doses in inhomogeneous radiation fields.
-In the context of IR, medical personnel are exposed to inhomogeneous radiation fields due to their proximity to the patient, which complicates the accurate assessment of individual doses.
+In the context of IR, medical staff are exposed to inhomogeneous radiation fields due to their proximity to the patient, which complicates the accurate assessment of individual doses.
 The reliability of current personal dosimetry methods is contingent upon the uniform distribution of radiation fields.
-However, this uniformity assumption does not hold in IR settings.
+However, this assumption of uniformity does not hold in IR settings.
 To address this issue, previous research has proposed the use of computational dosimetry systems [1], which are designed to monitor the locations and postures of all individuals involved in an IR procedure.
-In order to assess doses in these exposure scenarios, such systems are making use of currently available methods, that are mainly based on Monte-Carlo Simulations [2].
+In order to assess doses in these exposure scenarios, such systems make use of currently available methods, that are mainly based on Monte-Carlo Simulations [2].
 Nonetheless, the existing MCSs of radiation transport lack the necessary speed for real-time dose calculations, a limitation that also affects the software used for Virtual Reality (VR) or Augmented Reality (AR) training.
@@
 Furthermore, we have generated and made available a variety of datasets for the training of our models and analogous ones.
-The generation of our datasets was facilitated by the extension and utilization of the RadField3D [5] simulation software which itself is build upon the Geant4 [6] MCS framework to ensure physical correctness, as Geant4 is well established and validated for the targeted use-case.
+The generation of our datasets was facilitated by the extension and utilization of the RadField3D [5] simulation software which itself is built upon the Geant4 [6] MCS framework to ensure physical correctness, as Geant4 is well established and validated for the targeted use-case.
 Our primary innovation is the creation of three datasets of spatially resolved, three-dimensional fluence and spectra distributions for the training and evaluation of neural networks for the reconstruction of radiations fields for IR procedures.
-We accompany those with concrete neural network implementations and design recommendations based on ablation studies.
+We accompany these with concrete neural network implementations and design recommendations based on ablation studies.
 Additionally, we define several metrics useful for evaluating and comparing spatially resolved radiation field estimators.
@@
 Therefore, a set of well-established and tested general-purpose MCS toolkits for radiation transport already exists, such as Geant4 [6], MCNP [7], and EGSnrc [8].
-Since these toolkits are meant for general purpose applications ranging from astrophysics over nuclear reactor physics to dosimetry calculations, fine-tuning the MCS parameters for each specific use case is required.
-There are frameworks built upon these toolkits, such as GATE [9], which is based on Geant4, and it has already been used for assessing doses to medical staff in nuclear medicine by tailoring simulation models on the basis of real tracking of personnel movements [10].
+As these toolkits are meant for general purpose applications ranging from astrophysics over nuclear reactor physics to dosimetry calculations, fine-tuning the MCS parameters for each specific use case is required.
+There are frameworks built upon these toolkits, such as GATE [9], which is based on Geant4, and it has already been used for assessing doses to medical staff in nuclear medicine by tailoring simulation models based on the real tracking of personnel movements [10].
 However, as highlighted by Rondon et al., current MCS capabilities are still far from real-time as they require, at best, several tens of seconds if not minutes to be performed.
 This is particularly challenging in IR, as the calculation of radiation-fluence distributions in the space around the patient during interventions is computationally demanding and can require hours of simulations.
-However, real-time simulations capabilities are becoming more and more necessary for computational radiation protection dosimetry of staff, as proposed by prior research projects, such as the PODIUM project [1].
+However, real-time simulation capabilities are becoming increasingly necessary for computational radiation protection dosimetry of staff, as proposed by prior research projects, such as the PODIUM project [1].
 Besides the real-time dosimetry, the AR and VR applications for training interventional radiologists and associated personnel demand even faster execution speeds.
-Nevertheless, especially in AR and VR we can trade absolute accuracy for speed within reasonable bounds.
-Those applications were already presented various times of the last years, for example in 2012 by an c-arm training extension for VirtX [11] or by the perceptual study of Rainford et al.in 2023 [3].
-All of the already implemented systems would benefit from physically correct radiation fields to display during training, but currently those systems are only able to render radiation fields generated by simplified deterministic algorithms due to strict timing requirements.
-For those applications, achieving frame rates of at least $90-120\,\mathrm{fps}$ requires total rendering times per volume of approximately $8-11\,\mathrm{ms}$ to avoid perceptible visual latency [12].
+Nevertheless, especially in AR and VR, we can trade absolute accuracy for speed within reasonable bounds.
+These applications have already been presented at various times in the last years, for example in 2012 by a c-arm training extension for VirtX [11] or by the perceptual study of Rainford et al.in 2023 [3].
+All current implemented systems would benefit from the display of physically correct radiation fields during training, but at the moment these systems are only able to render radiation fields generated by simplified deterministic algorithms due to strict timing requirements.
+For these applications, achieving frame rates of at least $90$ to $120\,\mathrm{fps}$ requires total rendering times per volume of approximately $8$ to $11\,\mathrm{ms}$ to avoid perceptible visual latency [12].
 1.2 Computational Personal Dosimetry
 Real-world measurements in IR scenarios employing Active Personal Dosimeters have already revealed the complexity of contemporary radiation fields and high local dose rates that professionals are exposed to [13].
-Those measurements also implied that the definition of a representational location for wearing a dosimeter on the body of the medical staff (whether it is an APD or legal passive dosimeter) is not possible, and therefore, it is likely to both under- or overestimate the received doses by the staff.
+These measurements also implied that the definition of a representational location for wearing a personal dosemeter on the body of the medical staff is not possible, and therefore, it is likely to both under- or overestimate the received doses by the staff.
 In the course of the PODIUM project, the researchers developed a camera-based tracking system that was integrated with a MCS application based on the MCNP [7] toolkit.
@@
 To address this issue, recent studies have explored the prospect of enhancing its calculation speed.
-These efforts entail the replacement of the most computationally inefficient components of the radiation transport simulations, which is refereed to as a hybrid approach.
-Those hybrid methods can be composed of denoising algorithms.
+These efforts entail the replacement of the most computationally inefficient components of the radiation transport simulations, which is referred to as a hybrid approach.
+These hybrid methods can be composed of denoising algorithms.
 The most promising results in this field are delivered by deep learning methods, that use artificial neural networks for transforming the noisy results of an early aborted MCS to those of a fully executed MCS, but with shorter execution times [17, 18, 19].
-Although this approach has been the subject of more extensive research, it exhibits a significant disadvantage, as the MCS needs a certain minimum of statistics to be denoise-able which inevitably comes at the cost of a lower processing speed compared to a single neural network.
-Consequently, it does not meet the requirements for real-time applications especially in the context of VR and AR applications that we mentioned before.
+Although this approach has been the subject of more extensive research, it exhibits a significant disadvantage, as the MCS needs a certain minimum of statistics to be denoise-able which inevitably comes at the price of a lower processing speed compared to a single neural network.
+Consequently, it does not meet the requirements for real-time applications especially in the context of VR and AR applications previously mentioned.
 Other approaches focus on the complete replacement of the MCS algorithm, with artificial neural networks.
@@
 Concurrently, Ye et al.implemented a 3D- Convolutional Neural Network (CNN) for 3D reconstruction in a similar scenario [21].
-This 3D-CNN takes the unshielded radiation fields and shielding voxel geometry, convolutes those two channels down to a flat vector which is then rearranged to a 3D radiation field after the shielding process.
+This 3D-CNN takes the unshielded radiation fields and shielding voxel geometry, convolutes those two channels down to a flat vector which is then rearranged as a 3D radiation field after the shielding process.
 Another use case for FCNN is demonstrated in the research work of Yisheng et al.[22], where a FCNN is used to predict the spatially resolved photon flux in a room based on the data of a single detector at a known position.
@@
 2 Materials and Methods
-As there is a lack of public available and in this context usable datasets, our first contribution is the generation and provision of three datasets for training spatially resolved doserate predicting neural networks.
-For those three datasets, we implemented two distinct neural networks architectures, with additional variants for each, in order to analyze the influence of those architectures and specific hyperparameters on the quality of the predictions.
-For this purpose, we considered two basic architectures: A FCNN architecture inspired by Neural Radiance Fields [24] and a 3D-U-Net [25] architecture which is based on the CNN architecture, but with an extra transposed path for reconstructing all voxels at once.
+As there is a lack of publicly available, and in this context, usable datasets our first contribution is the generation and provision of three datasets for training spatially resolved dose rate predicting neural networks.
+For these three datasets, we implemented two distinct neural network architectures with additional variants for each, in order to analyze the influence of these architectures and specific hyperparameters on the quality of the predictions.
+For this purpose, we considered two basic architectures: an FCNN architecture inspired by Neural Radiance Fields [24] and a 3D-U-Net [25] architecture which is based on the CNN architecture, but with an extra transposed path for reconstructing all voxels at once.
 2.1 Datasets
@@
 The datasets were all generated using the RadField3D [5] dataset generator.
-The resulting datasets, consists of one radiation field per fixed set of parameters.
-Those fields are stored in the RadFiled3D file format and contain three channels.
-One for the photons of the direct X-ray beam, one for the resulting scatter fields and one containing the binary voxel occupation of the modeled geometry.
+The resulting datasets consist of one radiation field per fixed set of parameters.
+These fields are stored in the RadFiled3D file format and contain three channels.
+One for the photons of the direct X-ray beam, one for the resulting scattered fields and one containing the binary voxel occupation of the modeled geometry.
 Each of the first two channels contains the following information per voxel: The simulated volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$, the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ and the statistical error $\epsilon_{rel}$ as described in the research work about the MCS application RadField3D.
@@
 Each field has a resolution of 50×50×50 voxels with each voxel having an extent of $2\,\mathrm{cm}$ in every dimension resulting in a total field extent of $1\,\mathrm{m}$ per dimension.
-The varying parameter in this first dataset is the direction of the beam with $\phi,\theta\in[0,2\pi)$, while the X-ray tubes output is fixed to a radiation quality of H-100 according to ISO 4037-1 [26] and the beam shape is fixed to a cone beam with an opening angle of $\angle_{\mathrm{Beam}}=10\degree$.
-The distance of the X-ray tube relative to the origin is $2.5\,\mathrm{m}$.
-At the origin of the field, the headless torso of an male alderson RANDO phantom is placed to act as the scatter object for the radiation beam.
+The varying parameter in this first dataset is the direction of the beam with $\phi,\theta\in[0,2\pi)$, while the X-ray tubes output is fixed to a reference radiation quality of H-100 according to ISO 4037-1 [26] and the beam shape is fixed to a cone beam with an opening angle of $\angle_{\mathrm{Beam}}=10\degree$.
+The distance from the X-ray tube relative to the origin is $2.5\,\mathrm{m}$.
+At the origin of the field, the headless torso of a male Alderson RANDO phantom is placed to act as the scatter object for the radiation beam.
 This setting is derived from the validation measurements conducted for the RadField3D data generator [5], and, although simplified, it is considered realistic for modeling this first exposure scenario.
 DS-02:
-The second dataset contains $2156$ radiation fields and extends the previous one by the radiation source spectra as an varying parameter.
+The second dataset contains $2156$ radiation fields and extends the previous one by the radiation source spectra as a varying parameter.
 Here we used random, but realistic, spectra for modeling primary radiation in the context of IR.
-The range of the parameters was defined on the basis of empirical feedback from medical physicist and radiologists, mainly gained during the PODIUM project.
+The range of the parameters was defined based on empirical feedback from medical physicist and radiologists, mainly gained during the PODIUM project.
 Therefore, we defined an energy range for the tube output energy $E_{\mathrm{tube}}\in[40,125]\,\mathrm{keV}$.
 Further, we added aluminum with a random thickness $t_{\mathrm{Al}}\in[2.5,7.5]\,\mathrm{mm}$ and, optionally, copper, with a random-thickness $t_{\mathrm{Cu}}\in[0.0,0.9]\,\mathrm{mm}$ for simulating realistic filtering material used to improve the X-ray beam quality.
-As the anode material, we set tungsten, which is a standard reference material in IR imaging devices, and we used a random anode angle of $\angle_{\mathrm{Anode}}\in[8,12]\degree$.
+As the anode material, we set tungsten, which is a standard reference material in IR imaging devices, and used a random anode angle of $\angle_{\mathrm{Anode}}\in[8,12]\degree$.
 From that parameter space for the radiation source, we sampled sets of parameters and passed them to SpekPy [27] to generate reasonable X-ray spectra for the X-ray beams in this dataset.
@@
 In the third and most complex dataset, dynamic X-ray tube distancing is introduced.
-Therefore, a total of $3779$ distinct radiation fields was generated.
+A total of $3779$ distinct radiation fields were generated.
 For this dataset, we carried over the same base parameter space as described in the preceding datasets, and extended it by sampling the X-ray tube distance from the scenes isocenter as $d_{\mathrm{tube}}\in[35,75]\,\mathrm{cm}$.
@@
 In Tab.
-1, we have summarized the distribution of each input parameter inside the three datasets and additionally the distribution of the contained fluences by providing the dynamic ranges $\mathrm{DR}_{\mathrm{dB}}$ and the Gini coefficient $G_{\mathrm{Gini}}$, which are defined as:
-$\mathrm{DR}_{\mathrm{dB}}=10\log_{10}\Bigl(\frac{\max(\Phi^{3}_{\mathrm{\gamma}})}{\min(\Phi^{3}_{\mathrm{\gamma}})}\Bigr)$
+1, we have summarized the distribution of each input parameter inside the three datasets and additionally the distribution of the contained fluences by providing the dynamic ranges ${DR}_{\mathrm{dB}}$ and the Gini coefficient $G_{\mathrm{Gini}}$, which are defined as:
+${DR}_{\mathrm{dB}}=10\log_{10}\Bigl(\frac{\max(\Phi^{3}_{\mathrm{\gamma}})}{\min(\Phi^{3}_{\mathrm{\gamma}})}\Bigr)$
 $G_{\mathrm{Gini}}=\frac{1}{n^{2}\mu}\sum_{i=1}^{n}x_{i}(2i-n-1)$
-where $\mathrm{DR}_{\mathrm{dB}}$ is used to describe the actual ranges needed for reconstruction and $G_{\mathrm{Gini}}$ expresses the deviation of the present fluence distribution from a uniform distribution.
+where ${DR}_{\mathrm{dB}}$ is used to describe the actual ranges needed for reconstruction and $G_{\mathrm{Gini}}$ expresses the deviation of the present fluence distribution from a uniform distribution.
 For the calculation of $G_{\mathrm{Gini}}$, the fluences are expected to be sorted in ascending order.
@@
 Table 1: Statistics over the presented datasets including all varying input parameters.
-Dataset $\overline{\angle}_{\mathrm{Beam}}\pm\sigma$ $\overline{\mathrm{dist}}_{tube}\pm\sigma$ $\overline{\mathrm{kV}}\pm\sigma$ $\overline{\mathrm{kV_{p}}}\pm\sigma$ $\mathrm{DR}_{\mathrm{dB}}$ $G_{\mathrm{Gini}}$
-DS-01 $10.0\degree\pm 0.0$ $2.5\,\mathrm{m}\pm 0.0$ $80.7\,\mathrm{kV}\pm 0.0$ $100\,\mathrm{kV}\pm 0.0$ $43.9\,\mathrm{dB}$ $0.790$
-DS-02 $10.0\degree\pm 0.0$ $2.5\,\mathrm{m}\pm 0.0$ $48.7\,\mathrm{kV}\pm 9.3$ $82.1\,\mathrm{kV}\pm 24.6$ $53.5\,\mathrm{dB}$ $0.872$
-DS-03 - $0.6\,\mathrm{m}\pm 0.1$ $48.7\,\mathrm{kV}\pm 9.3$ $82.1\,\mathrm{kV}\pm 24.6$ $76.0\,\mathrm{dB}$ $0.991$
-By observing the differences between our three datasets, one can see, that by reducing the volume of the direct beam inside the radiation field volume, the dynamic range $\mathrm{DR}_{\mathrm{dB}}$ is also drastically increased.
-In the case of DS-03 when smaller opening angles of the beam and smaller distances of the X-ray tube from the isocenter are used, particle fluences $\Phi^{3}_{\mathrm{\gamma}}$ span over nearly $8$ orders of magnitudes, instead of the spanning of about $5$ orders of magnitudes for the other datasets.
-Concurrently, the distribution of fluences is more challenging, as the amount of voxels with fluences below $0.1\,\%$ of the maximum fluence rises by a factor of approximately $253$ from $0.3\,\%$ to $76\,\%$.
+Dataset $\overline{\angle}_{\mathrm{Beam}}\pm\sigma$ $\overline{dist}_{\mathrm{tube}}\pm\sigma$ $\overline{\mathrm{E}}\pm\sigma$ $\overline{\mathrm{keV_{p}}}\pm\sigma$ ${DR}_{\mathrm{dB}}$ $G_{\mathrm{Gini}}$
+DS-01 $10.0\degree\pm 0.0$ $2.5\,\mathrm{m}\pm 0.0$ $80.7\,\mathrm{keV}\pm 0.0$ $100.0\,\mathrm{keV}\pm 0.0$ $43.9\,\mathrm{dB}$ $0.790$
+DS-02 $10.0\degree\pm 0.0$ $2.5\,\mathrm{m}\pm 0.0$ $48.7\,\mathrm{keV}\pm 9.3$ $82.1\,\mathrm{keV}\pm 24.6$ $53.5\,\mathrm{dB}$ $0.872$
+DS-03 - $0.6\,\mathrm{m}\pm 0.1$ $48.7\,\mathrm{keV}\pm 9.3$ $82.1\,\mathrm{keV}\pm 24.6$ $76.0\,\mathrm{dB}$ $0.991$
+By observing the differences between our three datasets, one can see, that by reducing the volume of the direct beam inside the radiation field volume, the dynamic range ${DR}_{\mathrm{dB}}$ is drastically increased.
+In the case of DS-03 when smaller opening angles of the beam and smaller distances of the X-ray tube from the isocenter are used, particle fluences $\Phi^{3}_{\mathrm{\gamma}}$ span over nearly $8$ orders of magnitude, instead of the spanning of about $5$ orders of magnitude for the other datasets.
+Concurrently, the distribution of fluences is more challenging, as the amount of voxels with fluences below $0.1\,\%$ of the maximum fluence, rises by a factor of approximately $253$ from $0.3\,\%$ to $76\,\%$.
 2.2 Data normalization
-Figure 1: Distribution of relative voxel fluences $\Phi^{3}_{\mathrm{\gamma}}$ inside datasets DS-01 and DS-02 starting from 0.0 on the X-axis.
-Figure 2: Distribution of relative voxel fluences $\Phi^{3}_{\mathrm{\gamma}}$ inside datasets DS-01 and DS-02 starting from $10^{-1}$ on the X-axis.
-Figure 3: Distribution of relative voxel fluences $\Phi^{3}_{\mathrm{\gamma}}$ inside dataset DS-03 starting from 0.0 on the X-axis.
-Figure 4: Distribution of relative voxel fluences $\Phi^{3}_{\mathrm{\gamma}}$ inside dataset DS-03 starting from $10^{-1}$ on the X-axis.
+Figure 1: Distribution of relative voxel fluences inside datasets DS-01 and DS-02 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 0.0$ on the X-axis, exhibiting a strong imbalance towards $\Phi^{3}_{\mathrm{\gamma}}\leq 10^{-2}$.
+Figure 2: Distribution of relative voxel fluences inside datasets DS-01 and DS-02 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 10^{-1}$ on the X-axis, exhibiting a much weaker imbalance in that range.
+Figure 3: Distribution of relative voxel fluences inside dataset DS-03 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 0.0$ on the X-axis, exhibiting a strong imbalance towards $\Phi^{3}_{\mathrm{\gamma}}\leq 10^{-5}$.
+Figure 4: Distribution of relative voxel fluences inside dataset DS-03 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 10^{-1}$ on the X-axis, exhibiting a strong imbalance towards $\Phi^{3}_{\mathrm{\gamma}}\leq 20^{-1}$ and a small peak at $\Phi^{3}_{\mathrm{\gamma}}=1$.
 In principle, radiation fields inside the dataset are already normalized, as the photon spectra are normalized to probability distributions $p(E_{\mathrm{\gamma}})$ and the fluences are normalized per single primary photon from the simulation.
-However, due to the numerical nature of the datasets, after loading a single radiation field from a dataset, we performed an extra normalization step for the data to limit the occurring gradients and avoid numerical instability as well as introducing bias because of high total values.
-Thus this extra normalization step effectively mitigates the severe class imbalance in all our dataset towards low fluence voxels.
+However, due to the numerical nature of the datasets, after loading a single radiation field from a dataset, we performed an extra normalization step for the data to limit the occurring gradients and avoid numerical instability, as well as introducing bias because of high total values.
+Thus, this extra normalization step effectively mitigates the severe class imbalance in all our dataset towards low fluence voxels.
 This has been visualized in the global fluence histograms in Fig.
@@
 The histograms were generated using the MaxNorm function to transform the fluences to $[0,1]$.
-From both figures we can draw, that the majority of voxels has a fluence lower then $5\,\%$ of the maximum fluence, which makes sense, as the major part of the radiation field is indeed scatter radiation.
-However, this imbalance could potentially lead to a misreconstruction of the primary beam.
+From both figures we can draw that the majority of voxels have a fluence lower than $5\,\%$ of the maximum fluence, which makes sense, as the major part of the radiation field is indeed scattered radiation.
+However, this imbalance could potentially lead to a mis-reconstruction of the primary beam.
 Especially for DS-03 from Fig.
@@
 The idea of using alternative normalizations in the interval of $[0,1]$ is to create a more equal distribution of gradient strengths over the range of values to reconstruct.
-We pose it as a hypothesis, that this is of importance for our use case, as the dosimetry is regarding relative instead of absolute errors per voxel.
-Thus, it can be beneficial to elevate differences between small fluences and flatten the differences between higher fluences to give small deviations on small scales more impact.
+We hypothesize, that this is of importance for our use case, as the dosimetry is regarding relative instead of absolute errors per voxel.
+In general, it can be beneficial to elevate differences between small fluences and flatten the differences between higher fluences to give small deviations on small scales more impact.
 A classical solution for this issue is to use a transformation to the log-space, which is in general a reasonable choice for the following training of neural networks.
-Unfortunately, due to numerical instabilities, the steep ascend of the logarithm at very small values can become unstable under inversion and thus break the bijective property required for normalization.
-This instability can be observed for fluences of a magnitude lower than $10^{-5}$.
+Unfortunately, due to numerical instabilities, the steep ascent of the logarithm at very small values can become unstable under inversion and thus break the bijective property required for normalization.
+This instability can be observed for fluences of a magnitude lower than $10^{-8}$.
 As it can be seen from Tab.
-1, for DS-03 the interval of $10^{-7},10^{-8}]$ is of interest for the reconstruction as well.
-To mitigate this issue we added a scaling factor $\alpha$ prior to the log-transformation.
-We test two values for $\alpha$: $10^{3}$ and $10^{5}$.
-As an alternative way to encounter the high dynamic range, we doubled the range of our previously linearly spread out normalization from $\Phi^{3}_{\mathrm{\gamma}}\in[0,1]$ to $\Phi^{3}_{\mathrm{\gamma}}\in[-1,1]$ which additionally separates the low fluence voxels from the few high energy voxels by the sign.
+1, for DS-03 fluences of about $10^{-8}$ play a role for the reconstruction as well.
+To mitigate this issue, we added a scaling factor $\alpha$ prior to the log-transformation.
+We tested two values for $\alpha$: $10^{0}$ and $10^{3}$.
+As an alternative to encounter the high dynamic range, we doubled the range of our previously linearly spread-out normalization from $\Phi^{3}_{\mathrm{\gamma}}\in[0,1]$ to $\Phi^{3}_{\mathrm{\gamma}}\in[-1,1]$ which additionally separates the low fluence voxels from the few high energy voxels by the sign.
 We distinguish between those two variants of the MaxNorm by indicating them as $\mathrm{MaxNorm}_{0}^{1}$ and $\mathrm{MaxNorm}_{-1}^{+1}$.
 2.3 Architectures
-In order to access the effectivity of our neural network architectures, we implemented each architectures neural network model in three variants to handle the different provided parameters for each dataset.
+In order to access the effectivity of our neural network architectures, we implemented each architecture’s neural network model in three variants to handle the different provided parameters for each dataset.
 With this approach, for each architecture there is one model that can only handle X-ray tube rotations, one that can additionally handle varying spectra and one that can additionally handle varying tube distances.
-In general, our models are all estimating the relative topology of the spatial radiation field distribution conserving the relation of intensities between all voxels inside one specific radiation field.
+In general, our models all estimate the relative topology of the spatial radiation field distribution conserving the relation of intensities between all voxels inside one specific radiation field.
 This is important, as the actual fluence scales between different settings of the beam would introduce a bias towards high X-ray tube currents and thus harm the training of the neural networks.
-On the downside, in the real world application, this inherently requires our approach to run concurrently to at least one area dosimeter that reports doserates at a known point in space.
-As those devices, especially those with real-time measurement capabilities, are known to exhibit a certain energy and thus spectrum dependence, one needs to correct those devices accordingly.
-This is specifically of importance in our targeted use case, as we aim to predict the scatter field around the patient, where the spectrum is not homogeneous and difficultly predictable.
-Thus, we decided to train all our networks not of the actual target measurand, e.g., air kerma rate $\dot{K}_{\mathrm{air}}$ or $\mathbf{H}^{*}(10)$, but on the components of those measurands as they are provided by RadField3D, which are the volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$ and the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ composed of $32$ bins and $E_{max}=150\,\mathrm{keV}$.
-In fact, the concurrent measurements of an area dosimeter are only needed to correct the fluence.
-Therefore, one can potentially use the estimated spectra to correct an area dosimeter measurement in turn.
+On the downside, in a real-world application, this inherently requires our approach to run concurrently to at least one area dosemeter that reports dose rates at a known point in space.
+As these devices, especially those with real-time measurement capabilities, are known to exhibit a certain energy and thus spectrum dependence, one needs to correct them accordingly.
+This is specifically of importance in our targeted use case, as we aim to predict the scattered field around the patient, where the spectrum is not homogeneous and is difficult to predict.
+Thus, we decided to train all our networks not using the actual target measurand, e.g., air kerma rate $\dot{K}_{\mathrm{air}}$ free-in-air or $H^{*}(10)$, but on the components of those measurands as provided by RadField3D, which are the volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$ and the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ composed of $32$ bins and $E_{max}=150\,\mathrm{keV}$.
+In fact, the concurrent measurements of an area dosemeter are only needed to correct the fluence.
+Therefore, one can potentially use the estimated spectra to correct an area dosemeter measurement in turn.
 2.3.1 Fully-Connected-Voxelwise-Networks.
@@
 As we chose to not only predict a scalar dose rate per voxel, but the photon fluence $\Phi^{3}_{\mathrm{\gamma}}$ together with the normalized local photon spectrum $p(E_{\mathrm{\gamma}})$, this matches the NeRF architecture, where the reconstruction of visual density and color of a given point is separated as well.
-For the training process of our FCNNs, we used a cosine annealing [29] learning rate scheduler with $\mathrm{eta}_{min}=10^{-6}$ together with an Adam optimizer using weight decay [30] of $10^{-4}$ and $\beta_{1}=0.9\land\beta_{2}=0.99$.
+For the training process of our FCNNs, we used a cosine annealing [29] learning rate scheduler with $\mathrm{eta}_{\mathrm{min}}=10^{-6}$ together with an Adam optimizer using weight decay [30] of $10^{-4}$ and $\beta_{1}=0.9\land\beta_{2}=0.99$.
 Additionally, we preceded our main learning rate scheduler by a linear warmup phase that reaches our target learning rate after $1000$ steps.
@@
 5, where global parameters g in this case is the encoded direction vector.
-Our network begins by feeding the location vector through a frequency encoding using sine and cosine functions with different periodicities which was initially introduced by A.
+Our network begins by feeding the location vector through a encoder using frequency encoding using sine and cosine functions with different periodicities which were initially introduced by A.
 Vaswani et al.[31] for the transformer networks and reused for the original NeRF implementation.
-Even though, the direction vector could also be represented by a simple frequency encoding, we instead used a spherical harmonics based encoding as the cyclic nature of an rotation angle can better be represented by that encoding.
+Even though, the direction vector can also be represented by a simple frequency encoding, we instead used a spherical harmonics based encoding as the cyclic nature of a rotation angle can be better represented by that encoding.
 This was demonstrated for NeRF-like networks by InstantNGP of T.
 Müller et al.[32].
-For efficiency reasons, we use the spherical harmonics implementation as it is provided by the tiny-cuda-nn package [33] with version 2.0.
-For both encodings, we append the raw input to the encoded vector.
+For efficiency reasons, we used the spherical harmonics implementation as provided by the tiny-cuda-nn package [33] with version 2.0.
+For both encodings, we appended the raw input to the encoded vector.
 In contrast to the classical NeRF application, our network does not aim to learn the scatter objects 3D geometry with visual densities for volumetric rendering.
-Thus, we skip the pre-processing path where the model should optimize the voxel location representation towards the local density independently from the direction.
-We want the network to learn the implicit representation of the fluence distribution, which is similar to the visual density, but in our case is ultimately depended from the radiation direction.
-Thus, we extract the fluence as well as the spectrum from the last layer.
-In order build a robust backbone for solving the radiation field estimation under more complex parameter spaces, we do not just concatenate the feature vectors that describe the global beam conditions like the view direction in NeRF.
-Instead, we experimented with various feature fusing methods and finally compared the Feature-wise Linear Modulation (FiLM) [34] approach, which was originally used in CNNs, with a Gated Multimodal Unit (GMU) [35] approach and as the baseline, with the simple concatenation of feature vectors like it is done in classical NeRF networks.
-Ahead of fusing both feature vectors, the encoded X-ray tube direction vector is passed through the global encoding block, a Multilayer Perceptron (MLP) block that is similar to the MLP processing the encoded location vector with one hidden layer and a model width that is equal to the width of the main path of the network.
-The actual used hyperparameters for this and the extensions of this network, like frequencies of the frequency encoding, spherical harmonics degrees, model widths and feature fusing layers are reported in the results section as we determined them using hyperparameter tuning.
+Thus, we skipped the pre-processing path where the model should optimize the voxel location representation towards the local density independently from the direction.
+We want the network to learn the implicit representation of the fluence distribution, which is similar to the visual density, but in our case is ultimately dependent on the radiation direction.
+Thus, we extracted the fluence as well as the spectrum from the last layer.
+In order to build a robust backbone for solving the radiation field estimation under more complex parameter spaces, we did not just concatenate the feature vectors that describe the global beam conditions, such as the view direction in NeRF.
+Instead, we experimented with various feature fusing methods and finally compared the Feature-wise Linear Modulation (FiLM) [34] approach, which was originally used in CNNs, with a Gated Multimodal Unit (GMU) [35] approach and as the baseline, with the simple concatenation of feature vectors as is done in classical NeRF networks.
+Ahead of fusing both feature vectors, the encoded X-ray tube direction vector was passed through the global encoding block, a Multilayer Perceptron (MLP) block that is similar to the MLP processing the encoded location vector with one hidden layer and a model width that is equal to the width of the main path of the network.
+The actual hyperparameters used for this and the extensions of this network, such as the frequencies of the frequency encoding, spherical harmonics degrees, model widths and feature fusing layers are reported in the results section as we determined them using hyperparameter tuning.
 Through all layers of the main path of this network, except for the last ones, we use SiLU activation functions.
-Only for the output activation we use different activation functions.
-For the fluence, we use a different activation function depending on the range of the fluence normalization.
+Only in the case of the output activation did we use different activation functions.
+For the fluence, we used a different activation function depending on the range of the fluence normalization.
 For the case of a normalization in $[0,1]$, we used a sigmoid $\sigma$ activation function, while for a normalization in [1,1][-1,1], we used a gradient conserving clamping $\mathrm{clamp}_{\mathrm{GC}}$, with:
@@
 Each radiation field file contains the $150$ bin histogram of the X-ray tube output spectrum using a bin width of $1\,\mathrm{keV}$.
-In order to improve the radiation field distribution and include the reconstruction of the local spectra, we encode that spectrum into our network.
-First, we reduce the dimensionality drastically by bilinear resampling the bins of the histogram down to $64$ bins.
-The reduced histogram is then passed through a MLP with one hidden layer using a width of $32$ neurons and the SiLU activation function.
+In order to improve the radiation field distribution and include the reconstruction of the local spectra, we encoded that spectrum into our network.
+First, we reduced the dimensionality drastically by bilinear resampling the bins of the histogram down to $64$ bins.
+The reduced histogram was then passed through a MLP with one hidden layer using a width of $32$ neurons and the SiLU activation function.
 After the first linear layer, we implemented a layer-normalization to reshape the activation distribution.
-The two feature vectors of the X-ray tube direction and the X-ray tube output spectrum are then concatenated and combined by the global encoding block in Fig.
-5 that previously just processed the direction vector.
+The two feature vectors of the X-ray tube direction and the X-ray tube output spectrum were then concatenated and combined by the global encoding block in Fig.
+5 that previously processed the direction vector.
 SPERFNet was designed to be trained on the parameter space provided by DS-02.
 Parametric Beam Radiation Field Network (PBRFNet):
-This neural network targets the limitation of the pervious described variant, that is only able to learn a fixed, static beam distance.
+This neural network targets the limitation of the previous described variant, that is only able to learn a fixed, static beam distance.
 Therefore, this model adds an additional encoder to allow for the additional beam parameter introduced by DS-03 while keeping the same overall structure from SPERFNet.
-This requires one new scalar input parameter, that is the tube distance as the opening angle of the rectangular beam is directly depending on the distance.
-We observed, that simply passing this scalar raw is leading the neural network to ignore it.
+This requires one new scalar input parameter, that is the tube distance as the opening angle of the rectangular beam is directly dependent on the distance.
+We observed, that simply passing this scalar raw leads the neural network to ignore it.
 The previously used frequency encoding is not useful here either, as the input data is neither of high frequency nor periodical.
-Thus, we added a mini MLPs into the architecture, with has a width of $16$ neurons and in an optimal way encodes the distance information for the global parameters vector g.
+Thus, we added a mini MLPs into the architecture, which has a width of $16$ neurons and in an optimal way encodes the distance information for the global parameters vector g.
 PBRFNet was designed to be trained on the parameter space provided by DS-03.
@@
 Instead of predicting every single voxel by itself, which is inefficient when the aim is to always predict all voxels, especially for fine resolutions, one can use CNNs.
-Those networks are relying on using convolution kernels, with multiple channels, that sweep over the image space and by that encoding local information.
-Initially, those networks were only used for Computer Vision, but by transposing this operation, those networks can be used for generative purposes.
+These networks rely on using convolution kernels with multiple channels that sweep over the image space and by that, encodes local information.
+Initially, these networks were only used for Computer Vision, but by transposing this operation, these networks can be used for generative purposes.
 Beam2Scatter U-Net:
-Our Beam2Scatter (B2S) U-Net is in general a CNN, that encodes the raw radiation field of the direct beam prior to the interaction with the patient geometry and from that should estimate the scatter radiation field as it is induced by the patient.
+Our Beam2Scatter (B2S) U-Net is in general a CNN, that encodes the raw radiation field of the direct beam prior to the interaction with the patient geometry and from that estimates the scattered radiation field as it is induced by the patient.
 Therefore, its encoder-decoder architecture follows the 3D-U-Net description [25], which itself is just an adaption to the third dimension of the original U-Net, as it was initially used for biomedical image segmentation [36].
-We adopt this structure and give our network the fluence map of the direct beam as a one-channel input and let the network predict a $33$ dimensional output with the first $32$ channels for the spectra and the last channel for the fluence.
+We adopted this structure and gave our network the fluence map of the direct beam as a one-channel input and let the network predict a $33$ dimensional output with the first $32$ channels for the spectra and the last channel for the fluence.
 Another addition is, that we introduced FiLM [34] layers after each up-sampling step in the decoder path, to modulate the weights of the decoder depending on the encoded input spectrum of the X-ray tube, which is encoded the same way as in SPERFNet.
 This way we can inject that information into the U-Net structure.
-The networks outputs got the prior introduced activation functions applied: ${\mathrm{clamp}_{\mathrm{GC}}}_{min}^{max}$ and $\mathrm{norm}_{\mathrm{hist}}$.
-Thanks to the direct beam radiation distribution being the input of this network, it can be directly used with all three datasets.
+The networks outputs had the prior introduced activation functions applied: ${\mathrm{clamp}_{\mathrm{GC}}}_{min}^{max}$ and $\mathrm{norm}_{\mathrm{hist}}$.
+Thanks to the direct beam radiation distribution being the input of this network, it can be used directly with all three datasets.
 For DS-01, the spectra encoding and fusing by using FiLM layers was disabled.
@@
 Regardless, whether the model itself was capable of predicting partial or whole fields, the evaluation of the loss functions and the sub sequential update of the weights was performed in the same way.
-All models were trained for a maximum of $200$ epochs, which could be prematurely aborted, when the validation loss was not improving for the last $10$ epochs.
+All models were trained for a maximum of $200$ epochs, which could be prematurely aborted when the validation loss was not improving for the last $10$ epochs.
 For the training loop, we used the pytorch lightning [37] framework which implements the Cross-Validation policy with a final test using a dedicated validation and test dataset.
 The ratio of the split into training, validation and test dataset was $70\,\%$, $15\,\%$ and $15\,\%$.
-At first, a number of fields is loaded from the dataset that are transformed to pytorch tensors, preserving the channels and single layers, and as such they are batched together.
+At first, a number of fields were loaded from the dataset that are transformed to pytorch tensors, preserving the channels and single layers, and as such they were batched together.
 We used a physical batch size of $4$ full radiation fields.
-Multiple batches of those predicted and ground-truth fields are aggregated together to an effective batch size of $64$ for training.
-We applied data pre-processings to each field prior to the prediction.
-Those pre-processings included the optional joining of the direct beam and scattered field channels, but also normalizing the ground-truth and input tensors.
+Multiple batches of those predicted and ground-truth fields were aggregated together to an effective batch size of $64$ for training.
+We applied data pre-processing to each field prior to the prediction.
+This pre-processing included not only the option of joining the direct beam and scattered field channels, but also normalizing the ground-truth and input tensors.
 Finally, the loss functions for each of the neural networks normalized outputs (fluence and spectra) are calculated.
-Different from that, the metrics are calculated on the back-transformed data in the original data space.
-The initial learning rates where estimated per neural network by using the pytorch lighting implementation of a learning rate finder, which follows the descriptions of L.
+The metrics are calculated on the inverse-transformed data in the original data space.
+The initial learning rates were estimated per neural network by using the pytorch lighting implementation of a learning rate finder, which follows the descriptions of L.
 N.
@@
 For the FCNN models, which operate per voxel, we added an optional extra step to the training pipeline, that is typical for NeRF approaches.
-We refer to it as voxel location randomization, which is essentially a perturbation of the voxel position around its center with a maximum jitter of half the voxels dimension.
+We refer to it as voxel location randomization, which is essentially a perturbation of the voxel position around its center with a maximum jitter of half the voxel’s dimension.
 This was used as a data augmentation to stabilize training in NeRF variants like InstantNGP [32].
 2.4.1 Loss functions.
-Across all our models we used the same two loss functions, that we optimize concurrently: $\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}}$ the loss regarding the volumetric fluence $\Phi^{3}_{\mathrm{\gamma}}$ and $\pazocal{L}_{p(E_{\mathrm{\gamma}})}$ the loss regarding the local photon spectrum $p(E_{\mathrm{\gamma}})\in\mathbb{R}^{32}$ for $\sum_{i=0}^{31}p(E_{\mathrm{\gamma},i})=1$.
+Across all our models we used the same two loss functions, that we optimized concurrently: $\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}}$ the loss regarding the volumetric fluence $\Phi^{3}_{\mathrm{\gamma}}$ and $\pazocal{L}_{p(E_{\mathrm{\gamma}})}$ the loss regarding the local photon spectrum $p(E_{\mathrm{\gamma}})\in\mathbb{R}^{32}$ for $\sum_{i=0}^{31}p(E_{\mathrm{\gamma},i})=1$.
 $\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}}(t,p)$ $=\frac{1}{3}\Bigl(\mathrm{L_{1}}(t,p)+\mathrm{L_{2}}(t,p)+\bigl(1-\mathrm{SSIM}(t,p)\bigr)\Bigr),\$$\mathrm{with}\ t,p\in\mathbb{R}$
@@
 For the hyperparameter tuning, we used the Optuna [39] framework with version 4.5.
-For the tuning of each network we defined the search space for possible parameters as listed in Tab.
+For the tuning of each network, we defined the search space for possible parameters as listed in Tab.
 2.
-The tuning was in general performed on every model we discuss: B2S U-Net, SRBFNet, SPERFNet, PBRFNet.
+The tuning was, in general, performed on every model we discuss: B2S U-Net, SRBFNet, SPERFNet, PBRFNet.
 Anyway, not all of the parameters could be tested for each model.
@@
 feature fusing [Concatenation, FiLM, ResFiLM, GMU]
-normalizer $\mathrm{MaxNorm}_{0}^{1}$, $\mathrm{MaxNorm}_{-1}^{+1}$, $\mathrm{MaxLogNorm}_{10^{3}}$, $\mathrm{MaxLogNorm}_{10^{5}}$]
+normalizer $\mathrm{MaxNorm}_{0}^{1}$, $\mathrm{MaxNorm}_{-1}^{+1}$, $\mathrm{MaxLogNorm}_{10^{0}}$, $\mathrm{MaxLogNorm}_{10^{3}}$]
 random voxel location [true, false]
@@
 To assess the quality of the developed models, we need to analyze their performance on the test section of our datasets.
-As the distribution of voxel values inside our dataset is special for our use case and we have special requirements as well, e.g., we do not aim for perceptual accuracy as in most other reconstruction tasks, we first need to assess which metrics and which variants of metrics are applicable for neural network evaluation of this kind.
-In general, we measure our metrics over the air kerma rate fields $\dot{K}_{\mathrm{air}}$ of ground-truth and predicted fields using the following three metrics classes.
+As the distribution of voxel values inside our dataset are unique to our use case and we have special requirements as well, we did not aim for perceptual accuracy as in most other reconstruction tasks, we first needed to assess which metrics and which variants of metrics are applicable for neural network evaluation of this kind.
+In general, we measure our metrics over the air kerma rate free-in-air fields $\dot{K}_{\mathrm{air}}$ of ground-truth and predicted fields using the following three metrics classes.
 Symmetric Mean Absolute Percentage Error (SMAPE):
-The first and most obvious kind of metric we employed is the SMAPE which is providing the mean relative error across all voxels of the test dataset.
-As the dataset is strongly biased towards low voxel values we introduce this metric in two variants: $\mathrm{SMAPE}_{acc}^{x\,\%}$ for the metric applied over all voxels whose values exceed $\max(K_{air})*(1-\frac{x}{100})$ and $\mathrm{SMAPE}_{acc}^{scatter}$, which applies the metric on all voxels between $5\%$ and $0.5\%$ of $\max(K_{air})$.
-The SMAPE is defined in the following, where $\mathbf{P}$ and $\mathbf{T}$ are two sets of voxels, present in two corresponding $K_{air}$ fields.
+The first and most obvious kind of metric we employed is the SMAPE which provides the mean relative error across all voxels of the test dataset.
+As the dataset is strongly biased towards low voxel values we introduced this metric in two variants: $\mathrm{SMAPE}_{acc}^{x\,\%}$ for the metric applied over all voxels whose values exceed $\max(K_{\mathrm{air}})*(1-\frac{x}{100})$ and $\mathrm{SMAPE}_{acc}^{scatter}$, which applies the metric on all voxels between $5\,\%$ and $0.5\,\%$ of $\max(K_{\mathrm{air}})$.
+The SMAPE is defined in the following, where $\mathbf{P}$ and $\mathbf{T}$ are two sets of voxels, present in two corresponding $K_{\mathrm{air}}$ fields.
 $\mathbf{T}$ is the set of ground truth voxels and $\mathbf{P}$ is the set of predicted voxels:
@@
 Structural Similarity Index Measure (SSIM):
-This metric is a well established approach of measuring the similarity in terms of structure between a target and a prediction image or volume [40].
+This metric is a well established approach to measuring the similarity in terms of structure between a target and a prediction image or volume [40].
 This metric is commonly used for reconstruction tasks and expresses the structural similarity between target and prediction by an intuitive value between 0 and $1$, where $1$ indicates perfect accordance between both structures.
@@
 The GPR within the scope of a gamma evaluation is a common metric in the field of radiotherapy radiation transport simulation validation [41].
-It reflects the relative amount of voxels that pass a certain criterion within a given radiation field.
+It reflects the relative number of voxels that pass a certain criterion within a given radiation field.
 The metric variant we implemented follows the idea of the Maximum Allowed Dose Difference (MADD) as described by Jiang et al.[42].
 For our validation we used two different criteria.
-The strict criteria, which measures the amount of voxels with a MADD of $3\,\%$ using a spatial resolution of $4\,\mathrm{cm}$ and the more relaxed criterion, which measures the amount of voxels with a MADD of $10\,\%$ and a spatial resolution of $6\,\mathrm{cm}$.
-To calculate the GPR, the field is shifted by the allowed $\Delta_{d}$ in each dimension, respecting the extend of a single voxel.
+The strict criteria, which measures the number of voxels with a MADD of $3\,\%$ using a spatial resolution of $4\,\mathrm{cm}$ and the more relaxed criterion, which measures the number of voxels with a MADD of $10\,\%$ and a spatial resolution of $6\,\mathrm{cm}$.
+To calculate the GPR, the field is shifted by the allowed $\Delta_{d}$ in each dimension, respecting the extent of a single voxel.
 For $\Delta_{d}=4\,\mathrm{cm}$ this results in $2$ shifts per dimension for the case of a voxel size of $2\,\mathrm{cm}$.
@@
 $\gamma^{\Delta_{d}}_{\Delta_{\dot{D}}}(\vec{x})=\min_{\vec{r}\in\mathbf{V}_{\Delta_{d}}(\vec{x})}\Biggl(\sqrt{\Bigl(\frac{\|\vec{x}-\vec{r}\|}{\Delta_{d}}\Bigr)^{2}+\Bigl(\frac{\dot{D}_{p}(\vec{x})-\dot{D}_{t}(\vec{r})}{\Delta_{\dot{D}}}\Bigr)^{2}}\Biggr),\mathbf{V}_{\Delta_{d}}(\vec{x})=\{\vec{v}\,|\,\|\vec{v}-\vec{x}\|<={\Delta_{d}}\land\vec{v}\in\mathbf{X}\}$
-with $\dot{D}_{t}$ being the target doserate in $K_{air}$ at a given point and $\dot{D}_{p}$ being the predicted one.
+with $\dot{D}_{t}$ being the target dose rate in $K_{\mathrm{air}}$ at a given point and $\dot{D}_{p}$ being the predicted one.
 The reported metric $\mathrm{GPR}_{p\,\%}^{d\,\mathrm{cm}}$ follows from that as:
@@
 Spectrum Accuracy:
-As the spectra represent a special kind of information with additional usage aside the direct dosimetry aspect, we evaluate them separately.
+As the spectra represent a special kind of information that can be used for purposes other than direct dosimetry, we evaluate them separately.
 Therefore, we use the Intersection over Union (IoU), where we calculate the ratio between the combined area of ground truth and prediction spectra and the overlapped area between them.
@@
 3.
-In there, each metric value belongs to a network, trained, tested and optimized in terms of hyperparameters on one of our datasets.
-The first dataset DS-01, was used to train our base FCNN architecture called SRBF.
+Therein, each metric value belongs to a network, trained, tested and optimized in terms of hyperparameters on one of our datasets.
+The first dataset DS-01, was used to train our base FCNN architecture called SRBFNet.
 We compared the results of that network against the implementation of the U-Net that we introduced previously as Beam-to-Scatter (B2S) U-Net.
@@
 On DS-01, one can easily see, that the FCNN architecture outperforms the U-Net architecture.
-Even though, the U-Net also learns the structure of the radiation field, the edges between primary beam and scattered radiation are not reproduced as sharp as the FCNN does.
+Even though, the U-Net also learns the structure of the radiation field, the edges between primary beam and scattered radiation are not reproduced as sharply as the FCNN does.
 This observation can be reproduced for DS-02.
@@
 Furthermore, we can conclude, that training our FCNN architecture on a bigger dataset using more degrees of freedom, namely the X-ray tube spectrum, actually improves the $\mathrm{SMAPE}_{acc}$, even though the spatial resolution got slightly, but not notably, worse.
-For DS-03 and PBRFNet, we observed, that using models with a higher neuron count by increasing the width of our model beyond $192$, e.g., $256$ or even $384$, the reconstruction capability is increasing by up to $1.5\,\%$ regarding $\mathrm{SMAPE}_{acc}^{scatter}$.
+For DS-03 and PBRFNet, we observed, that using models with a higher neuron count by increasing the width of our model per layer beyond $192$ neurons, e.g., $256$ or even $384$, the reconstruction capability increases by up to $1.5\,\%$ regarding $\mathrm{SMAPE}_{acc}^{scatter}$.
 A similar increase of the SSIM on the other hand side, and thus the overall shape of the radiation field, is invariant to a change of the model width beyond $192$ across all our datasets and models.
@@
 4.
-From those results, we can conclude, that for the fluence prediction, the same data range and activation function as used for NeRF-like approaches is beneficial, which is the use of sigmoid $\sigma$ together with a linear maximum normalization between $[0,1]$.
+From those results, we can conclude, that for the fluence prediction, the same data range and activation function as used for NeRF-like approaches is beneficial, which is using sigmoid $\sigma$ together with a linear maximum normalization between $[0,1]$.
 Moreover, when additionally reviewing the table of Tab.
@@
 In that table, we concentrate on the differences regarding $\mathrm{SMAPE}_{acc}^{scatter}$, as this metric is reflecting the accuracy on the most interesting part of the predicted volume for the computational dosimetry.
-That is, because the staff during an IR procedure will be mainly monitored in the $\mathrm{SMAPE}_{acc}^{scatter}$ volume.
+That is because the staff, during an IR procedure, will be mainly monitored in the $\mathrm{SMAPE}_{acc}^{scatter}$ volume.
 From Tab.
-5, one can also draw, that the use of FiLM layers did not only achieve the best accuracies, but also the shortest inference durations compared to GMU or Residual-FiLM.
+5, one can also draw the conclusion, that the use of FiLM layers did not only achieve the best accuracies, but also the shortest inference durations compared to GMU or Residual-FiLM.
 Moreover, we demonstrated, that the radiation transport simulation can be learned by the same deep, but narrow structure that is used to learn light transfer as in the use case of NVS.
@@
 Nevertheless, an inference time of about $20\,\mathrm{ms}$ is suitable for interactive applications.
-After all, with the provided FCNN models, state-of-the-art inferences times, that are suitable for the VR- and AR-application, should be achievable by implementing the network as CUDA or OpenCL kernels in the future, like it was already demonstrated by InstantNGP.
+After all, with the provided FCNN models, state-of-the-art inferences times, that are suitable for the VR- and AR-application, should be achievable by implementing the network as CUDA or OpenCL kernels in the future, as has already been demonstrated by InstantNGP.
 5 Conclusion

### v3

@@
-Estimating Spatially Resolved Radiation Fields Using Neural Networks
+Learning-Based Estimation of Spatially Resolved Scatter Radiation Fields in Interventional Radiology
 Abstract
-We present an in-depth analysis on how to build and train neural networks to estimate the spatial distribution of scattered radiation fields for radiation protection dosimetry in medical radiation fields, such as those found in interventional radiology and cardiology.
-We present three different synthetically generated datasets with increasing complexity for training, using a Monte-Carlo Simulation application based on Geant4.
+We present three variants of a lightweight, fully connected artificial neural network, suited for interactive estimation of three-dimensional, spatially resolved volumes of scattered radiation fields and a corresponding training pipeline for radiation protection dosimetry in medical radiation fields, such as those found in interventional radiology and cardiology.
+Accompanying, we present three different synthetically generated datasets with increasing complexity for training, generated using RadField3D, a Monte Carlo simulation application based on Geant4.
+As the primary scatter object, we employed the torso of a male Alderson RANDO phantom.
 On those datasets, we evaluate convolutional and fully connected architectures of neural networks to demonstrate which design decisions work well for reconstructing the fluence and spectra distributions over the spatial domain of such radiation fields.
 All our datasets, as well as our training pipeline, are published as open source in separate repositories.
-Code: https://github.com/Centrasis/radfield3d-nn
+To evaluate the presented neural networks, we define and assess several metrics.
+Across these measures, the model variants demonstrate good spatial agreement between predicted and ground-truth radiation fields, particularly within specific regions of interest within the radiation field.
+Of particular relevance for potential application in out-of-field dosimetry is the SMAPE of the scatter radiation field, which represents the most challenging metric and was consistently above $84\,\%$.
 keywords:
-Interventional Radiology, Monte-Carlo Simulation, FCNN, U-Net, Dataset, X-ray
+Interventional Radiology, Monte Carlo Simulation, FCNN, U-Net, Dataset, X-ray
 articletype: Paper
@@
 MCS
-Monte-Carlo Simulation
+Monte Carlo Simulation
 DSR
-Dose-Structure Report
+Dose Structured Report
 CT
@@
 Gated Multimodal Unit
+SiLU
+Sigmoid Linear Unit
 ICG
@@
 To address this issue, previous research has proposed the use of computational dosimetry systems [21], which are designed to monitor the locations and postures of all individuals involved in an IR procedure.
-In order to assess doses in these exposure scenarios, such systems make use of currently available methods, that are mainly based on Monte-Carlo Simulations [2].
+In order to assess doses in these exposure scenarios, such systems make use of currently available methods, that are mainly based on Monte Carlo Simulations [6].
 Nonetheless, the existing MCSs of radiation transport lack the necessary speed for real-time dose calculations, a limitation that also affects the software used for Virtual Reality (VR) or Augmented Reality (AR) training.
@@
 To advance radiation protection, it is necessary to develop accelerated methods for the simulation of radiation transport.
-In this research work, we explore the surrogate model approach for MCS acceleration, whereby we learn spatially resolved radiation field distributions from a Monte-Carlo Simulation.
+In this research work, we explore the surrogate model approach for MCS acceleration, whereby we learn spatially resolved radiation field distributions from a Monte Carlo Simulation.
 Subsequently, the neural networks serve as efficient models for radiation-field reconstruction, providing a real-time capable radiation-transport simulation method.
@@
 As these toolkits are meant for general purpose applications ranging from astrophysics over nuclear reactor physics to dosimetry calculations, fine-tuning the MCS parameters for each specific use case is required.
-There are frameworks built upon these toolkits, such as GATE [9], which is based on Geant4, and it has already been used for assessing doses to medical staff in nuclear medicine by tailoring simulation models based on the real tracking of personnel movements [10].
-However, as highlighted by Rondon et al., current MCS capabilities are still far from real-time as they require, at best, several tens of seconds if not minutes to be performed.
+There are frameworks built upon these toolkits, such as GATE [13], which is based on Geant4, and it has already been used for assessing doses to medical staff in nuclear medicine by tailoring simulation models based on the real tracking of personnel movements as demonstrated by Rondón et al.
+[30].
+However, as highlighted by Rondón et al., current MCS capabilities are still far from real-time as they require, at best, several tens of seconds if not minutes to be performed.
 This is particularly challenging in IR, as the calculation of radiation-fluence distributions in the space around the patient during interventions is computationally demanding and can require hours of simulations.
@@
 Here, a 3D-UNet is used to transform a CT-Scout-Scan to the surrounding radiation field around the patient induced by the CT-scanner.
+This work aims to contribute to the corpus of existing surrogate model approaches by focusing on predicting dynamic radiation fields using lightweight FCNNs.
+This enables the prediction of partial radiation fields within the region of interest, thereby supporting real-time capable radiation field estimation and, in theory, also allowing for higher resolutions of radiation fields.
+In contrast to prior approaches, our approach additionally provides the local energy spectra, making it possible to use those networks for correcting the energy dependence of real world dosemeters.
+In the following, we present FCNNs inspired by the architecture of the networks used for Neural Radiance Fields [18].
+As these networks model geometry implicitly, and our datasets do not incorporate variations in patient geometry, a direct comparison with the work of Arias et al.
+[35] is neither meaningful nor feasible, given that their approach explicitly encodes patient geometry.
 2 Materials and Methods
@@
 For these three datasets, we implemented two distinct neural network architectures with additional variants for each, in order to analyze the influence of these architectures and specific hyperparameters on the quality of the predictions.
-For this purpose, we considered two basic architectures: an FCNN architecture inspired by Neural Radiance Fields [24] and a 3D-U-Net [25] architecture which is based on the CNN architecture, but with an extra transposed path for reconstructing all voxels at once.
+For this purpose, we considered two basic architectures: an FCNN architecture inspired by NeRFs [18] and a 3D-U-Net [7] architecture which is based on the CNN architecture, but with an extra transposed path for reconstructing all voxels at once.
 2.1 Datasets
@@
 The datasets were all generated using the RadField3D [15] dataset generator.
+Per field, the generation took about $20\,\mathrm{min}$ on our AMD Epyc 9474F 48-core CPU.
 The resulting datasets consist of one radiation field per fixed set of parameters.
@@
 One for the photons of the direct X-ray beam, one for the resulting scattered fields and one containing the binary voxel occupation of the modeled geometry.
-Each of the first two channels contains the following information per voxel: The simulated volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$, the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ and the statistical error $\epsilon_{rel}$ as described in the research work about the MCS application RadField3D.
-Additionally, all input parameters, such as the X-ray tubes output spectra or the tubes position and direction, are stored per field.
+Each of the first two channels contains the following information per voxel: The simulated volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$ (fluence per cubic volume of a voxel), the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ and the statistical error $\epsilon_{rel}$ as described in the research work about the MCS application RadField3D.
+Additionally, all input parameters, such as the X-ray tube output spectra or the tubes position and direction, are stored per field.
 DS-01:
@@
 Each field has a resolution of 50×50×50 voxels with each voxel having an extent of $2\,\mathrm{cm}$ in every dimension resulting in a total field extent of $1\,\mathrm{m}$ per dimension.
-The varying parameter in this first dataset is the direction of the beam with $\phi,\theta\in[0,2\pi)$, while the X-ray tubes output is fixed to a reference radiation quality of H-100 according to ISO 4037-1 [26] and the beam shape is fixed to a cone beam with an opening angle of $\angle_{\mathrm{Beam}}=10\degree$.
+The varying parameter in this first dataset is the direction of the beam with $\phi,\theta\in[0,2\pi)$, while the X-ray tube output is fixed to a reference radiation quality of H-100 according to ISO 4037-1 [9] and the beam shape is fixed to a cone beam with an opening angle of $\angle_{\mathrm{Beam}}=10\degree$.
 The distance from the X-ray tube relative to the origin is $2.5\,\mathrm{m}$.
-At the origin of the field, the headless torso of a male Alderson RANDO phantom is placed to act as the scatter object for the radiation beam.
+At the origin of the field, the headless torso of a male Alderson RANDO phantom [27] is placed to act as the scatter object for the radiation beam.
 This setting is derived from the validation measurements conducted for the RadField3D data generator [15], and, although simplified, it is considered realistic for modeling this first exposure scenario.
@@
 The range of the parameters was defined based on empirical feedback from medical physicist and radiologists, mainly gained during the PODIUM project.
-Therefore, we defined an energy range for the tube output energy $E_{\mathrm{tube}}\in[40,125]\,\mathrm{keV}$.
+Therefore, we defined an energy range for the tube output peak energy $E_{\mathrm{tube}}\in[40,125]\,\mathrm{keV}$.
 Further, we added aluminum with a random thickness $t_{\mathrm{Al}}\in[2.5,7.5]\,\mathrm{mm}$ and, optionally, copper, with a random-thickness $t_{\mathrm{Cu}}\in[0.0,0.9]\,\mathrm{mm}$ for simulating realistic filtering material used to improve the X-ray beam quality.
@@
 A total of $3779$ distinct radiation fields were generated.
-For this dataset, we carried over the same base parameter space as described in the preceding datasets, and extended it by sampling the X-ray tube distance from the scenes isocenter as $d_{\mathrm{tube}}\in[35,75]\,\mathrm{cm}$.
-In order to make the dataset even more realistic in representing typical real-world scenarios, the X-ray beam was collimated to a rectangular shape, with fixed dimensions of $40\,\mathrm{cm}\times 30\,\mathrm{cm}$ in the plane passing through the origin.
+For this dataset, we mostly carried over the same base parameter space as described in the preceding datasets, but in contrast to the previous datasets, we sampled the X-ray tube distance from the scenes isocenter as $d_{\mathrm{tube}}\in[35,75]\,\mathrm{cm}$.
+In order to make the dataset even more realistic in representing typical real-world scenarios, the X-ray beam was collimated to a rectangular shape, with fixed dimensions of $40\,\mathrm{cm}\times 30\,\mathrm{cm}$ in the plane passing through the origin and thus intersecting the phantom.
+Table 1: Listing of the statistical errors, with respect to the fluence, from the scattered radiation fields separated by dataset.
+$\mathrm{mean}(error)$ refers to the mean over all voxels in the dataset, while $\mathrm{median}(error)$ refers to the mean over the median statistical error of each file inside the dataset.
+Dataset $\mathrm{mean}(error)$ $\mathrm{median}(error)$
+DS-01 $9.7\,\%$ $0.03\,\%$
+DS-02 $9.9\,\%$ $0.03\,\%$
+DS-03 $9.9\,\%$ $0.03\,\%$
+RadField3D uses the variation of the relative voxel fluences during the simulation process to measure the statistical error per voxel.
+Those statistical errors from the scattered radiation field layer are evaluated continuously to determine the current overall statistical error.
+Once $90\,\%$ of those voxels exhibit a statistical error equal to or less than $10\,\%$, the simulation is aborted.
+This behavior is reflected by Tab.
+1 where we show the mean and median statistical errors present in our previously described datasets, as calculated by RadField3D.
+An example of the spatial distribution of relative air kerma rates in air $\dot{K}_{\mathrm{air}}$ can bee seen in Fig.
+1.
+The two projections were generated from a random field taken from DS-03 by excluding all doserates below $0.5\,\%$ of the maximum doserate for improved visual understanding.
+Figure 1: Two perspective projections of a random radiation field from DS-03: One projection along the X-axis (left) and one along Y-axis (right).
+The spatially resolved relative doserate $\dot{K}_{\mathrm{air}}$ (air kerma in air) is color coded from transparent ($0.5\,\%$) over blue ($0.5\,\%$) to yellow ($100\,\%$).
+The used scatter geometry (Male Alderson RANDO phantom torso) is colored in red.
 In Tab.
-1, we have summarized the distribution of each input parameter inside the three datasets and additionally the distribution of the contained fluences by providing the dynamic ranges ${DR}_{\mathrm{dB}}$ and the Gini coefficient $G_{\mathrm{Gini}}$, which are defined as:
-${DR}_{\mathrm{dB}}=10\log_{10}\Bigl(\frac{\max(\Phi^{3}_{\mathrm{\gamma}})}{\min(\Phi^{3}_{\mathrm{\gamma}})}\Bigr)$
+2, we have summarized the distribution of each input parameter inside the three datasets and additionally the distribution of the contained fluences by providing the dynamic ranges ${DR}_{\mathrm{dB}}$ and the Gini coefficient $G_{\mathrm{Gini}}$, which are defined as:
+${DR}_{\mathrm{dB}}=10\log_{10}\Bigl(\frac{\max(\Phi^{3}_{\mathrm{\gamma}}\penalty 10000\ )}{\min(\Phi^{3}_{\mathrm{\gamma}}\penalty 10000\ )}\Bigr)$
 $G_{\mathrm{Gini}}=\frac{1}{n^{2}\mu}\sum_{i=1}^{n}x_{i}(2i-n-1)$
 where ${DR}_{\mathrm{dB}}$ is used to describe the actual ranges needed for reconstruction and $G_{\mathrm{Gini}}$ expresses the deviation of the present fluence distribution from a uniform distribution.
-For the calculation of $G_{\mathrm{Gini}}$, the fluences are expected to be sorted in ascending order.
-A Gini coefficient of $G_{\mathrm{Gini}}=0$ indicates, that the dataset is uniformly distributed, while $G_{\mathrm{Gini}}=1$ indicates, that the dataset contains only one value or category.
+For the calculation of $G_{\mathrm{Gini}}$, the fluences are expected to be sorted in ascending order , where $\mu$ denotes the mean fluence and $n$ the total count of voxels across the dataset.
+Finally, a Gini coefficient of $G_{\mathrm{Gini}}=0$ indicates that the dataset is uniformly distributed, while $G_{\mathrm{Gini}}=1$ indicates that all fluences of the dataset are accumulated in one single bin of the fluence histogram.
 Table 2: Statistics over the presented datasets including all varying input parameters.
@@
 Concurrently, the distribution of fluences is more challenging, as the amount of voxels with fluences below $0.1\,\%$ of the maximum fluence, rises by a factor of approximately $253$ from $0.3\,\%$ to $76\,\%$.
-2.2 Data normalization
-Figure 1: Distribution of relative voxel fluences inside datasets DS-01 and DS-02 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 0.0$ on the X-axis, exhibiting a strong imbalance towards $\Phi^{3}_{\mathrm{\gamma}}\leq 10^{-2}$.
-Figure 2: Distribution of relative voxel fluences inside datasets DS-01 and DS-02 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 10^{-1}$ on the X-axis, exhibiting a much weaker imbalance in that range.
-Figure 3: Distribution of relative voxel fluences inside dataset DS-03 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 0.0$ on the X-axis, exhibiting a strong imbalance towards $\Phi^{3}_{\mathrm{\gamma}}\leq 10^{-5}$.
-Figure 4: Distribution of relative voxel fluences inside dataset DS-03 for range $\Phi^{3}_{\mathrm{\gamma}}\geq 10^{-1}$ on the X-axis, exhibiting a strong imbalance towards $\Phi^{3}_{\mathrm{\gamma}}\leq 20^{-1}$ and a small peak at $\Phi^{3}_{\mathrm{\gamma}}=1$.
-In principle, radiation fields inside the dataset are already normalized, as the photon spectra are normalized to probability distributions $p(E_{\mathrm{\gamma}})$ and the fluences are normalized per single primary photon from the simulation.
-However, due to the numerical nature of the datasets, after loading a single radiation field from a dataset, we performed an extra normalization step for the data to limit the occurring gradients and avoid numerical instability, as well as introducing bias because of high total values.
-Thus, this extra normalization step effectively mitigates the severe class imbalance in all our dataset towards low fluence voxels.
-This has been visualized in the global fluence histograms in Fig.
-4 for DS-01 and DS-02 and in Fig.
-4 for DS-03.
-The histograms were generated using the MaxNorm function to transform the fluences to $[0,1]$.
-From both figures we can draw that the majority of voxels have a fluence lower than $5\,\%$ of the maximum fluence, which makes sense, as the major part of the radiation field is indeed scattered radiation.
-However, this imbalance could potentially lead to a mis-reconstruction of the primary beam.
-Especially for DS-03 from Fig.
-4, we can see that the class imbalance is also present inside the primary beam component.
-Not as severely expressed, datasets DS-01 and DS-02 exhibit an unequal distribution of fluences as well, as visible from Fig.
-4.
-Therefore, we examine the influence of different normalization functions regarding the reconstruction capabilities, to find the best fit to reconstruct the high and low fluence regions with sufficient results.
-$\mathrm{MaxNorm}(x)$ $=\frac{x}{\max(\mathbf{X})},$ $\quad x\in\mathbf{X}\subset\mathbb{R}^{+}$
-$\Rightarrow\mathrm{MaxNorm}:\mathbb{R}^{+}\rightarrow[0,1]$
-$\mathrm{MaxLogNorm}_{\alpha}(x)$ $=\frac{\ln(1+\alpha\,x)}{\max_{y\in\mathbf{X}}(\ln(1+\alpha\,y))},$ $\quad x\in\mathbf{X}\subset\mathbb{R}^{+}$
-$\Rightarrow\mathrm{MaxLogNorm}:\mathbb{R}^{+}\rightarrow[0,1]$
-The idea of using alternative normalizations in the interval of $[0,1]$ is to create a more equal distribution of gradient strengths over the range of values to reconstruct.
-We hypothesize, that this is of importance for our use case, as the dosimetry is regarding relative instead of absolute errors per voxel.
-In general, it can be beneficial to elevate differences between small fluences and flatten the differences between higher fluences to give small deviations on small scales more impact.
-A classical solution for this issue is to use a transformation to the log-space, which is in general a reasonable choice for the following training of neural networks.
-Unfortunately, due to numerical instabilities, the steep ascent of the logarithm at very small values can become unstable under inversion and thus break the bijective property required for normalization.
-This instability can be observed for fluences of a magnitude lower than $10^{-8}$.
-As it can be seen from Tab.
-1, for DS-03 fluences of about $10^{-8}$ play a role for the reconstruction as well.
-To mitigate this issue, we added a scaling factor $\alpha$ prior to the log-transformation.
-We tested two values for $\alpha$: $10^{0}$ and $10^{3}$.
-As an alternative to encounter the high dynamic range, we doubled the range of our previously linearly spread-out normalization from $\Phi^{3}_{\mathrm{\gamma}}\in[0,1]$ to $\Phi^{3}_{\mathrm{\gamma}}\in[-1,1]$ which additionally separates the low fluence voxels from the few high energy voxels by the sign.
-We distinguish between those two variants of the MaxNorm by indicating them as $\mathrm{MaxNorm}_{0}^{1}$ and $\mathrm{MaxNorm}_{-1}^{+1}$.
+Overall, for all three datasets, the dataset generator RadField3D ensures, that $90\,\%$ of all voxels do not exceed a statistical error of $10\,\%$ with respect to fluence.
+Data normalization:
+All datasets used are provided normalized by RadField3D.
+This applies both to the voxelwise photon energy spectra, which are probability distributions with unit integral, and to the voxelwise fluences, which are normalized with respect to the number primary particles of the simulation.
+This effectively scales all fluence values relative to the maximum, ensuring that all fluences fall within the $[0..1]$ interval.
 2.2 Architectures
@@
 This is specifically of importance in our targeted use case, as we aim to predict the scattered field around the patient, where the spectrum is not homogeneous and is difficult to predict.
-Thus, we decided to train all our networks not using the actual target measurand, e.g., air kerma rate $\dot{K}_{\mathrm{air}}$ free-in-air or $H^{*}(10)$, but on the components of those measurands as provided by RadField3D, which are the volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$ and the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ composed of $32$ bins and $E_{max}=150\,\mathrm{keV}$.
+Thus, we decided to train all our networks not using the actual target measurand, e.g., air kerma rate $\dot{K}_{\mathrm{air}}$ free-in-air or ambient dose equivalent $H^{*}(10)$, but on the components of those measurands as provided by RadField3D, which are the volumetric photon fluence $\Phi^{3}_{\mathrm{\gamma}}$ and the volumetric photon energy distribution $p(E_{\mathrm{\gamma}})$ composed of $32$ bins and $E_{max}=150\,\mathrm{keV}$.
 In fact, the concurrent measurements of an area dosemeter are only needed to correct the fluence.
-Therefore, one can potentially use the estimated spectra to correct an area dosemeter measurement in turn.
+Therefore, one can potentially use the estimated spectra to correct an area dosemeter measurement for its energy and angle dependence in turn.
 2.2.1 Fully-Connected-Voxelwise-Networks.
@@
 The overall structure is shown in Fig.
-5, where global parameters g in this case is the encoded direction vector.
-Our network begins by feeding the location vector through a encoder using frequency encoding using sine and cosine functions with different periodicities which were initially introduced by A.
+2, where global parameters g in this case is the encoded direction vector.
+Our network begins by feeding the location vector through an encoder using frequency encoding composed of sine and cosine functions with different periodicities which were initially introduced by A.
 Vaswani et al.[34] for the transformer networks and reused for the original NeRF implementation.
@@
 In order to build a robust backbone for solving the radiation field estimation under more complex parameter spaces, we did not just concatenate the feature vectors that describe the global beam conditions, such as the view direction in NeRF.
-Instead, we experimented with various feature fusing methods and finally compared the Feature-wise Linear Modulation (FiLM) [34] approach, which was originally used in CNNs, with a Gated Multimodal Unit (GMU) [35] approach and as the baseline, with the simple concatenation of feature vectors as is done in classical NeRF networks.
-Ahead of fusing both feature vectors, the encoded X-ray tube direction vector was passed through the global encoding block, a Multilayer Perceptron (MLP) block that is similar to the MLP processing the encoded location vector with one hidden layer and a model width that is equal to the width of the main path of the network.
-The actual hyperparameters used for this and the extensions of this network, such as the frequencies of the frequency encoding, spherical harmonics degrees, model widths and feature fusing layers are reported in the results section as we determined them using hyperparameter tuning.
-Through all layers of the main path of this network, except for the last ones, we use SiLU activation functions.
-Only in the case of the output activation did we use different activation functions.
-For the fluence, we used a different activation function depending on the range of the fluence normalization.
-For the case of a normalization in $[0,1]$, we used a sigmoid $\sigma$ activation function, while for a normalization in [1,1][-1,1], we used a gradient conserving clamping $\mathrm{clamp}_{\mathrm{GC}}$, with:
+Instead, we experimented with various feature fusion methods and finally compared the Feature-wise Linear Modulation (FiLM) [24] approach, which was originally used in CNNs, with a Gated Multimodal Unit (GMU) [4] approach and as the baseline, with the simple concatenation of feature vectors as is done in classical NeRF networks.
+Ahead of fusion both feature vectors, the encoded X-ray tube direction vector was passed through the global encoding block, a Multilayer Perceptron (MLP) block that is similar to the MLP processing the encoded location vector with one hidden layer and a model width that is equal to the width of the main path of the network.
+The actual hyperparameters used for this and the extensions of this network, such as the frequencies of the frequency encoding, spherical harmonics degrees, model widths and feature fusion layers are reported in the results section as we determined them using hyperparameter tuning.
+Through all layers of the main path of this network, except for the last ones, we use Sigmoid Linear Unit (SiLU) activation functions.
+For the output layer, we used two different activation functions.
+For the fluence, we used the sigmoid activation, to keep the output between 0 and 1:
 $\sigma(x)=\frac{1}{1+\mathrm{exp}(-x)}$
-${\mathrm{clamp}_{\mathrm{GC}}}_{min}^{max}(x)=x+(\mathrm{clamp}_{min}^{max}(x)-x)$
-For the gradient conserving clamping, the second term of the sum is executed detached from pytorchs autograd graph which results in a clamping that still has the original gradients attached.
 For the local spectra, we use a simple histogram normalization $\mathrm{norm}_{\mathrm{hist}}$ as the activation function, to guarantee normalized spectra for the losses.
@@
 This way we can inject that information into the U-Net structure.
-The networks outputs had the prior introduced activation functions applied: ${\mathrm{clamp}_{\mathrm{GC}}}_{min}^{max}$ and $\mathrm{norm}_{\mathrm{hist}}$.
+The networks outputs had the prior introduced activation functions applied: sigmoid and $\mathrm{norm}_{\mathrm{hist}}$.
 Thanks to the direct beam radiation distribution being the input of this network, it can be used directly with all three datasets.
-For DS-01, the spectra encoding and fusing by using FiLM layers was disabled.
+For DS-01, the spectra encoding and fusion by using FiLM layers was disabled.
 2.3 Training pipeline
@@
 All models were trained for a maximum of $200$ epochs, which could be prematurely aborted when the validation loss was not improving for the last $10$ epochs.
-For the training loop, we used the pytorch lightning [37] framework which implements the Cross-Validation policy with a final test using a dedicated validation and test dataset.
+For the training loop, we used the pytorch lightning [8] framework in version 2.5.6 which implements the Cross-Validation policy with a final test using a dedicated validation and test dataset.
 The ratio of the split into training, validation and test dataset was $70\,\%$, $15\,\%$ and $15\,\%$.
@@
 This was used as a data augmentation to stabilize training in NeRF variants like InstantNGP [19].
-2.4.1 Loss functions.
-Across all our models we used the same two loss functions, that we optimized concurrently: $\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}}$ the loss regarding the volumetric fluence $\Phi^{3}_{\mathrm{\gamma}}$ and $\pazocal{L}_{p(E_{\mathrm{\gamma}})}$ the loss regarding the local photon spectrum $p(E_{\mathrm{\gamma}})\in\mathbb{R}^{32}$ for $\sum_{i=0}^{31}p(E_{\mathrm{\gamma},i})=1$.
-$\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}}(t,p)$ $=\frac{1}{3}\Bigl(\mathrm{L_{1}}(t,p)+\mathrm{L_{2}}(t,p)+\bigl(1-\mathrm{SSIM}(t,p)\bigr)\Bigr),\$$\mathrm{with}\ t,p\in\mathbb{R}$
+2.3.1 Loss functions
+Across all our models we used the same two loss functions, that we optimized concurrently: $\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}\penalty 10000\ }$ the loss regarding the volumetric fluence $\Phi^{3}_{\mathrm{\gamma}}$ and $\pazocal{L}_{p(E_{\mathrm{\gamma}})}$ the loss regarding the local photon spectrum $p(E_{\mathrm{\gamma}})\in\mathbb{R}^{32}$ for $\sum_{i=0}^{31}p(E_{\mathrm{\gamma},i})=1$.
+$\pazocal{L}_{\Phi^{3}_{\mathrm{\gamma}}\penalty 10000\ }(t,p)$ $=\frac{1}{3}\Bigl(\mathrm{L_{1}}(t,p)+\mathrm{L_{2}}(t,p)+\bigl(1-\mathrm{SSIM}(t,p)\bigr)\Bigr),\$$\mathrm{with}\ t,p\in\mathbb{R}$
 […]
@@
 With $t$ being the ground truth (target) and $p$ being the prediction of the neural network.
+A quick exploration of the parameter space had lead to the presented ratio between Wasserstein and $L_{1}$-Loss and was beneficial for achieving satisfactory results.
 Here, $L_{1}$ and $L_{2}$ refer to the L1- and L2-Norms using mean aggregation to produce a scalar loss for higher dimensional inputs.
-2.4.2 Hyperparameter tuning.
+2.3.2 Hyperparameter tuning
 For the hyperparameter tuning, we used the Optuna [3] framework with version 4.5.
 For the tuning of each network, we defined the search space for possible parameters as listed in Tab.
-2.
+3.
 The tuning was, in general, performed on every model we discuss: B2S U-Net, SRBFNet, SPERFNet, PBRFNet.
@@
 direction encoding harmonics [4, 6, 8]
-feature fusing [Concatenation, FiLM, ResFiLM, GMU]
-normalizer $\mathrm{MaxNorm}_{0}^{1}$, $\mathrm{MaxNorm}_{-1}^{+1}$, $\mathrm{MaxLogNorm}_{10^{0}}$, $\mathrm{MaxLogNorm}_{10^{3}}$]
+feature fusion [Concatenation, FiLM, ResFiLM, GMU]
 random voxel location [true, false]
@@
 This metric is commonly used for reconstruction tasks and expresses the structural similarity between target and prediction by an intuitive value between 0 and $1$, where $1$ indicates perfect accordance between both structures.
+$\mathrm{SSIM}(p,t)=\frac{(2\mu_{p}\mu_{t}+C_{1})(2\sigma_{pt}+C_{2})}{(\mu_{p}^{2}+\mu_{t}^{2}+C1)(\sigma_{p}^{2}+\sigma_{t}^{2}+C_{2})}$
+with $\mu_{x}$ being the mean of the window $x$, $\sigma_{x}$ being the variance of the window $x$, and $\sigma_{xy}$ being the covariance matrix of the two windows $x$ and $y$.
+The $\mathrm{SSIM}(p,t)$ is calculated for small windows $p$ and $t$ of the compared volumes $\mathbf{P}$ and $\mathbf{T}$, and the global SSIM is calculated by the weighted sum of all windows.
+$C_{1}$ and $C_{2}$ are calculated from the dynamic range of the compared volumes as stabilization constants to prevent terms from degrading to zero.
 Gamma Passing Ratio (GPR):
@@
 $\mathbf{T}$ is the set of ground truth voxels and $\mathbf{P}$ is the set of predicted voxels:
-$\mathrm{Spec}_{acc}=\frac{\sum_{i=1}^{n}{\min(P_{i},T_{i})}}{\sum_{i=1}^{n}{P_{i}+T_{i}-\min(P_{i},T_{i})}}$
+$\mathrm{Spec}_{acc}=\frac{\sum_{i=1}^{n}{\min(P_{i},T_{i})}}{\sum_{i=1}^{n}{\bigl(P_{i}+T_{i}-\min(P_{i},T_{i})\bigr)}}$
 3 Results
-Table 3: Reported reconstruction metrics for each network in its optimal configuration of hyperparameters. All metrics were calculated from the test dataset section, grouped by used dataset.
-Dataset Model $\mathrm{SMAPE}_{acc}^{90\,\%}$ $\mathrm{SMAPE}_{acc}^{scatter}$ SSIM $\mathrm{GPR}_{3\,\%}^{6\,\mathrm{cm}}$ $\mathrm{GPR}_{10\,\%}^{4\,\mathrm{cm}}$ $\mathrm{Spec}_{acc}$
-DS-01 SRBF $\mathbf{96.4}\,\%$ $\mathbf{95.5}\,\%$ $\mathbf{0.961}$ $\mathbf{99.84\,\%}$ $\mathbf{99.98}\,\%$ $\mathbf{84.2}\,\%$
-B2S U-Net $87.1\,\%$ $64.0\,\%$ $0.868$ $84.9\,\%$ $99.0\%$ $72.5\,\%$
-DS-02 SRBF $90.0\,\%$ $83.3\,\%$ $0.923$ $6.8\,\% 99$$9.8\,\%$ $59.5\,\%$
-SPERF $\mathbf{97.3}\,\%$ $\mathbf{96.2}\,\%$ $\mathbf{0.952}$ $\mathbf{99.7\,\%}$ $\mathbf{99.9}\,\%$ $\mathbf{86.2}\,\%$
-B2S U-Net $88.0\,\%$ $55.5\,\%$ $0.844$ $72.7\,\%$ $98.1\%$ $77.1\,\%$
-DS-03 PBRF $\mathbf{84.4}\,\%$ $\mathbf{84.8}\,\%$ $\mathbf{0.902}$ $\mathbf{87.2}\,\%$ $\mathbf{94.4}\,\%$ $\mathbf{86.7}\,\%$
-B2S U-Net $19.5\,\%$ $19.4\,\%$ $0.802$ $21.8\,\%$ $30.5\,\%$ $14.2\,\%$
 In general, the qualities of the developed models by this research work are condensed in Tab.
-3.
+4.
 Therein, each metric value belongs to a network, trained, tested and optimized in terms of hyperparameters on one of our datasets.
@@
 We report the best metric values our networks achieved after we optimized the models hyperparameters by using Optuna as mentioned in the previous section.
-The specific found hyperparameters for each network, are listed in Tab.
-4.
+The specific found hyperparameters for each network are listed in Tab.
+5.
+Table 4: Reported reconstruction metrics for each network in its optimal configuration of hyperparameters.
+All metrics were calculated from the test dataset section, grouped by used dataset.
+For the training of the models a NVIDIA H100 GPU was used.
+Dataset Model $\mathrm{SMAPE}_{acc}^{90\,\%}$ $\mathrm{SMAPE}_{acc}^{scatter}$ SSIM $\mathrm{GPR}_{3\,\%}^{6\,\mathrm{cm}}$ $\mathrm{GPR}_{10\,\%}^{4\,\mathrm{cm}}$ $\mathrm{Spec}_{acc}$
+DS-01 SRBF $\mathbf{96.4}\,\%$ $\mathbf{95.5}\,\%$ $\mathbf{0.961}$ $\mathbf{99.8\,\%}$ $\mathbf{99.98}\,\%$ $\mathbf{84.2}\,\%$
+B2S U-Net $87.1\,\%$ $64.0\,\%$ $0.868$ $84.9\,\%$ $99.0\%$ $72.5\,\%$
+DS-02 SRBF $90.0\,\%$ $83.3\,\%$ $0.923$ $6.8\,\% 99$$9.8\,\%$ $59.5\,\%$
+SPERF $\mathbf{97.1}\,\%$ $\mathbf{96.2}\,\%$ $\mathbf{0.952}$ $\mathbf{99.7\,\%}$ $\mathbf{99.9}\,\%$ $\mathbf{86.2}\,\%$
+B2S U-Net $88.0\,\%$ $55.5\,\%$ $0.844$ $72.7\,\%$ $98.1\%$ $77.1\,\%$
+DS-03 PBRF $\mathbf{84.4}\,\%$ $\mathbf{84.8}\,\%$ $\mathbf{0.902}$ $\mathbf{87.2}\,\%$ $\mathbf{94.4}\,\%$ $\mathbf{86.7}\,\%$
+B2S U-Net $19.5\,\%$ $19.4\,\%$ $0.802$ $21.8\,\%$ $30.5\,\%$ $14.2\,\%$
 On DS-01, one can easily see, that the FCNN architecture outperforms the U-Net architecture.
-Even though, the U-Net also learns the structure of the radiation field, the edges between primary beam and scattered radiation are not reproduced as sharply as the FCNN does.
+Even though, the U-Net also learns the structure of the radiation field, as reflected by the SSIM scores in Tab.
+4, the edges between primary beam and scattered radiation are not reproduced as sharply as the FCNN does.
+This can be concluded from the relation between SMAPE accuracy on both regions and the SSIM scores, which drops less than the accuracy.
 This observation can be reproduced for DS-02.
@@
 The effectivity of our spectra encoding can be seen by comparing SRBFNet against SPERFNet, where we get a nominal difference of $12.9\,\%$ for $\mathrm{SMAPE}_{acc}^{scatter}$ just from the missing spectra information.
-Furthermore, we can conclude, that training our FCNN architecture on a bigger dataset using more degrees of freedom, namely the X-ray tube spectrum, actually improves the $\mathrm{SMAPE}_{acc}$, even though the spatial resolution got slightly, but not notably, worse.
+Furthermore, when comparing SRBF on DS-01 with SPERF on DS-02 in Tab.
+4, we can conclude, that training our FCNN architecture on a bigger dataset using more degrees of freedom, namely the X-ray tube spectrum, actually improves the $\mathrm{SMAPE}_{acc}$, even though the spatial resolution , according to the drop in the SSIM score, got slightly, but not notably, worse.
 For DS-03 and PBRFNet, we observed, that using models with a higher neuron count by increasing the width of our model per layer beyond $192$ neurons, e.g., $256$ or even $384$, the reconstruction capability increases by up to $1.5\,\%$ regarding $\mathrm{SMAPE}_{acc}^{scatter}$.
@@
 As the benefit of using broader models is not significant enough to sacrifice performance for it, we limited the hyperparameter selection to models with a maximum model width of $192$.
+From Tab.
+6, we can see, that the use of GMU feature fusion is doubling the inference durations and decreases the learning capabilities consistently over all datasets and FCNN based models.
+The use of FiLM feature fusion, on the other hand, is faster and often consistently beneficial for the SMAPE accuracy metric over the scattered radiation field region.
+In general, an inference duration for a whole radiation field estimation of about $20\,\mathrm{ms}$ was achieved by all our FCNN based models, when used with concatenation or FiLM based feature fusion.
+Please note, that the corresponding B2S-U-Net implementation achieved a mean inference duration of $7.5\,\mathrm{ms}\pm 2.0\,\mathrm{ms}$ on the same datasets and hardware using FiLM conditioning in the center layer.
 Table 5: Used model configurations, where L denotes the number of frequencies used for the Fourier feature encoding of the voxel location, $l_{max}$ denotes the maximum degree used for the encoding of the radiation direction by spherical harmonics and $\mathrm{dim}(\mathbf{g}_{\mathrm{spec}})$ indicates the dimensionality of the encoded input spectrum.
 Model width refers to the number of neurons in one full layer of the MLP main path.
-Dataset Model Model Width L $l_{max}$ $\mathrm{dim}(\mathbf{g}_{\mathrm{spec}})$ Norm Conditioning
-DS-01 SRBF 192 10 4 - $\mathrm{MaxNorm}_{0}^{1}$ FiLM
-B2S U-Net - - - - $\mathrm{MaxNorm}_{0}^{1}$ -
-DS-02 SRBF 192 10 4 - $\mathrm{MaxNorm}_{0}^{1}$ FiLM
-SPERF 192 10 4 16 $\mathrm{MaxNorm}_{0}^{1}$ FiLM
-B2S U-Net - - - - $\mathrm{MaxNorm}_{0}^{1}$ FiLM
-DS-03 PBRF 192 14 4 16 $\mathrm{MaxNorm}_{0}^{1}$ FiLM
-B2S U-Net - - - - $\mathrm{MaxNorm}_{0}^{1}$ FiLM
-Table 5: Inference durations per field ($50^{3}$ voxels) of each model, regarding a specific feature fusing method, using non-compiled pytorch models and a NVIDIA RTX 3090 GPU.
+Dataset Model Model Width L $l_{max}$ $\mathrm{dim}(\mathbf{g}_{\mathrm{spec}})$ Conditioning
+DS-01 SRBF 192 10 4 - FiLM
+B2S U-Net - - - - -
+DS-02 SRBF 192 10 4 - FiLM
+SPERF 192 10 4 16 FiLM
+B2S U-Net - - - - FiLM
+DS-03 PBRF 192 14 4 16 FiLM
+B2S U-Net - - - - FiLM
+Table 6: Inference durations per field ($50^{3}$ voxels) of each model, regarding a specific feature fusion method, using non-compiled pytorch models and a NVIDIA RTX 4090 GPU in a Intel i9-13900K system with $128\,\mathrm{GB}$ of RAM installed.
 All models were using the following configuration with a model width of $192$ and encoding parameters $l_{max}=4\land\mathrm{L}=10$ for SRBFNet and SPERFNet, with one exception for PBRFNet, which was using L = $14$.
-Model Feature fusing Inference duration per field $\mathrm{SMAPE}_{acc}^{scatter}$
-SRBF Concat $20.86\,\mathrm{ms}\ \pm\ 2.81\,\mathrm{ms}$ $94.3\,\%$
-FiLM $20.55\,\mathrm{ms}\ \pm\ 1.87\,\mathrm{ms}$ $\mathbf{95.5}\,\%$
-ResFiLM $24.46\,\mathrm{ms}\ \pm\ 1.82\,\mathrm{ms}$ $94.4\,\%$
-GMU $41.00\,\mathrm{ms}\ \pm\ 2.55\,\mathrm{ms}$ $88.7\,\%$
-SPERF Concat $21.76\,\mathrm{ms}\ \pm\ 2.57\,\mathrm{ms}$ $95.2\,\%$
-FiLM $21.76\,\mathrm{ms}\ \pm\ 2.57\,\mathrm{ms}$ $\mathbf{96.2}\,\%$
-ResFiLM $24.71\,\mathrm{ms}\ \pm\ 1.63\,\mathrm{ms}$ $95.4\,\%$
-GMU $41.71\,\mathrm{ms}\ \pm\ 2.51\,\mathrm{ms}$ $93.7\,\%$
-PBRF Concat $21.03\,\mathrm{ms}\ \pm\ 1.31\,\mathrm{ms}$ $79.2\,\%$
-FiLM $22.17\,\mathrm{ms}\ \pm\ 2.54\,\mathrm{ms}$ $\mathbf{84.8}\,\%$
-ResFiLM $24.32\,\mathrm{ms}\ \pm\ 1.88\,\mathrm{ms}$ $67.8\,\%$
-GMU $40.43\,\mathrm{ms}\ \pm\ 2.42\,\mathrm{ms}$ $40.9\,\%$
+Model Feature fusion Inference duration per field $\mathrm{SMAPE}_{acc}^{scatter}$
+SRBF Concat $20.7\,\mathrm{ms}\ \pm\ 2.8\,\mathrm{ms}$ $94.3\,\%$
+FiLM $20.6\,\mathrm{ms}\ \pm\ 1.9\,\mathrm{ms}$ $\mathbf{95.5}\,\%$
+ResFiLM $24.5\,\mathrm{ms}\ \pm\ 1.8\,\mathrm{ms}$ $94.4\,\%$
+GMU $41.0\,\mathrm{ms}\ \pm\ 2.6\,\mathrm{ms}$ $88.7\,\%$
+SPERF Concat $21.8\,\mathrm{ms}\ \pm\ 2.6\,\mathrm{ms}$ $95.2\,\%$
+FiLM $21.8\,\mathrm{ms}\ \pm\ 2.6\,\mathrm{ms}$ $\mathbf{96.2}\,\%$
+ResFiLM $24.7\,\mathrm{ms}\ \pm\ 1.6\,\mathrm{ms}$ $95.4\,\%$
+GMU $41.7\,\mathrm{ms}\ \pm\ 2.5\,\mathrm{ms}$ $93.7\,\%$
+PBRF Concat $21.0\,\mathrm{ms}\ \pm\ 1.3\,\mathrm{ms}$ $79.2\,\%$
+FiLM $22.2\,\mathrm{ms}\ \pm\ 2.5\,\mathrm{ms}$ $\mathbf{84.8}\,\%$
+ResFiLM $24.3\,\mathrm{ms}\ \pm\ 1.9\,\mathrm{ms}$ $67.8\,\%$
+GMU $40.4\,\mathrm{ms}\ \pm\ 2.4\,\mathrm{ms}$ $40.9\,\%$
 4 Discussion
 The configuration of the best performing models during the hyperparameter tuning are listed in Tab.
-4.
-From those results, we can conclude, that for the fluence prediction, the same data range and activation function as used for NeRF-like approaches is beneficial, which is using sigmoid $\sigma$ together with a linear maximum normalization between $[0,1]$.
-Moreover, when additionally reviewing the table of Tab.
-3, we are convinced, that using a model width of $192$ and a location encoding with frequencies within a range of $10$ to at max $14$ is sufficient for the presented use case.
+5.
+When additionally reviewing the table of Tab.
+4, we are convinced, that using a model width of $192$ and a location encoding with frequencies within a range of $10$ to at max $14$ is sufficient for the presented use case.
 Consistently over all datasets, the angular resolution of the beam needed to solve this problem was found to be low enough, that a harmonics count of $4$ for the spherical harmonics encoding of the beam direction is sufficient.
-Further, two plain FiLM layers for feature fusing, one at an early and one at a late position in the network, was proven to be most effective for this use case compared to using GMU layers, Residual-FiLM layers and feature concatenation.
+Further, two plain FiLM layers for feature fusion, one at an early and one at a late position in the network, was proven to be most effective for this use case compared to using GMU layers, Residual-FiLM layers and feature concatenation.
 Even though, the impact of the FiLM layers on the inference duration is measurable, it is small compared to the concatenation that is commonly used in the field of Novel View Synthesis, as it can be seen in Tab.
-5.
+6.
 In that table, we concentrate on the differences regarding $\mathrm{SMAPE}_{acc}^{scatter}$, as this metric is reflecting the accuracy on the most interesting part of the predicted volume for the computational dosimetry.
@@
 From Tab.
-5, one can also draw the conclusion, that the use of FiLM layers did not only achieve the best accuracies, but also the shortest inference durations compared to GMU or Residual-FiLM.
+6, one can also draw the conclusion, that the use of FiLM layers did not only achieve the best accuracies, but also the shortest inference durations compared to GMU or Residual-FiLM.
 Moreover, we demonstrated, that the radiation transport simulation can be learned by the same deep, but narrow structure that is used to learn light transfer as in the use case of NVS.
@@
 Nevertheless, an inference time of about $20\,\mathrm{ms}$ is suitable for interactive applications.
-After all, with the provided FCNN models, state-of-the-art inferences times, that are suitable for the VR- and AR-application, should be achievable by implementing the network as CUDA or OpenCL kernels in the future, as has already been demonstrated by InstantNGP.
+After all, with the provided FCNN models, state-of-the-art inferences times, that are suitable for the VR- and AR-application, should be achievable by implementing the network as CUDA or OpenCL kernels in the future, as has already been demonstrated by InstantNGP [19].
 5 Conclusion
@@
 Additionally, we have implemented and compared different variants of fully connected and convolutional neural networks, following the architectural design of U-Nets and NeRF-like architectures.
-Thereby, we have proven NeRF-like architectures to be efficient and capable of being used for radiation protection purposes and training of medical staff in the context of IR procedures.
+Our results indicate that NeRF-like architectures are effective and suitable for radiation protection applications, and are therefore likely applicable to the training of medical staff in the context of IR procedures.
 We have also pointed out, which design decisions are effective in this special context.
 Our presented datasets together with the presented findings on the architectural decisions for neural reconstruction networks create a profound foundation for enabling further research on this topic, e.g., extending our or similar approaches by uncertainty handling in neural networks or incorporation of dynamic scatter geometry.
+6 Future work
+We leave it as future work to implement the presented FCNN using fused CUDA kernels to improve their performance.
+In doing so, it is expected that the networks will achieve inference times suitable for AR/VR rendering, as we have presented lightweight NeRF-like architectures with two FiLM layers.
+We demonstrated that these FiLM layers only add minimal overhead.
+In general, this approach was already demonstrated for NeRF variants like InstantNGP [19].
+Further, we empathize the importance of public available datasets for training and comparing neural networks for dosimetry.
+Especially, more datasets with increasing realism are needed by adding beam collimation that is independent from the tubes distance from the isocenter or by adding various patient shapes and poses to the dataset generation process.
 Data availability